A Bibliography on Case-Based Reasoning

Friedrich Gebhardt GMD FIT-KI

Sankt Augustin D-53754 Germany

Tel: 02241-14-2660 FAX: 02241-14-2384

E-mail: friedrich.gebhardt@gmd.de (Internet)

30 January 2019 Version 1.09

Abstract

This bibliography records publications in the field of case-based reasoning, and related areas.

Title word cross-reference

(ML)^{2} [82, 103].

21st [57]. 2D [113]. 2nd [31, 100, 71].

56-3 [282].

’93 [321, 318]. ’94 [322, 324]. 95 [273].

A.S.A. [195]. A4 [106, 144, 145, 323]. AAAO [228]. abstract [235]. access [198]. achieved [172]. acquire [259]. Acquisition [54, 115, 242, 79]. action [19]. active [228, 231]. Adaptation [228, 283, 236, 179, 249, 217, 307, 308]. adaptive [241]. adaptiven [89]. Advances [324, 70]. advisory [170]. against [98]. Age [72, 5]. agent [19]. AgentEx [236]. agents [236]. Ähnlichkeit [101]. Ähnlichkeiten [97]. ähnlichkeitsbasierte [163]. Ähnlichkeitsurteilen [111]. AI [169, 95]. AI-Workbench [95]. aid [78, 96, 256]. aided [241, 243]. aids [188]. aktiver [153]. Algenmassenentwicklungen [272]. algorithm [63]. Allgemeine [48]. allgemeiner [227]. Analogical [237, 61, 176, 286, 47]. analogy [238, 286, 39]. Analysing [102]. Analysis [262, 294, 15, 66, 139, 114, 221, 71]. analytic [281]. Anciens [1]. andere [30]. annotated [17]. ANOPLA [252]. Anpassung [153]. Anpassungswissens [288]. Ansatz [230, 141]. Ansätze [67, 43]. Anwendung [267]. Anwendungen [69, 309]. Application [134, 31, 168, 266]. Applications [311, 152, 280]. Applying [254]. Approach [54, 107, 161, 171, 176, 179, 189, 320, 224, 130]. Approaches [314, 167]. aptitude [168]. Arbeitsplatz [117]. ARCADE [283]. architects [302]. architectural [231, 246, 195, 198, 256]. Architecture [20, 85, 243, 187, 96, 226]. arguments [299]. ARMILLA [35]. arrangement [122]. Art [49]. artifacts [278]. Artificial [100, 322, 320]. ASK [24]. ASLIB [8]. AspecT [269]. Aspekte [270]. Assessing [257]. Assessment [233, 232, 148, 208, 274]. assignment [109]. assistance [231, 283]. assistant [195, 276]. assisted [266]. associative [197]. atdex [88]. attention [29]. attribute [197, 212]. attribute-value [197, 212]. August [322, 320]. Ausführbare [125]. Automated [36]. Automatic [246, 247]. automatische [265, 253]. autonomer [153]. autonomous [228]. autonomy [99].

BABYLON [50, 95]. BABYLON-Wissensbasis [50]. Background [24]. Band [129]. Bankprojekt [135]. BankXX [299]. base [198]. Based [229, 54, 291, 115, 293, 311, 161, 273, 62, 167, 280, 281, 168, 131, 317, 92, 171, 175, 176, 133, 177, 178, 179, 180, 244, 181, 316, 187, 78, 96, 188, 52, 99, 81, 199, 315, 324, 254, 146, 202, 203, 108, 205, 206, 148, 319, 114, 150, 214, 85, 152, 297, 157, 158, 217, 88, 222, 123, 164, 275, 165, 321, 166, 279, 75]. Basic [262, 287]. Bauen [323, 106, 144]. Bauentwurf [153]. Bauentwurfsprozeß [155]. Bausteine [218, 227]. bauten [12]. Bauwesen [53]. be [220]. Beach [315]. Behandlung [43, 97]. behaviour [181, 6, 119]. Benutzung [248]. Betrieb [106, 144, 323]. betrieblicher [117]. between [169, 197, 212]. Beyond [18]. bibliography [17, 214]. Bilderkennung [120]. Boeing [282]. Bound [64, 63]. Bound-Verfahren [64].

Branch [64, 63]. Bridging [169]. Buch [77]. Building [235, 132, 99, 36, 126, 285, 240, 243, 145, 201, 255, 275, 166, 279]. Buildings [59, 145]. built [170].

C [113]. C-string [113]. CAD

[45, 220, 221]. CAD-plans [220, 221].

CADET [123]. CADRE [146]. Call [309]. Call-Center-Anwendungen [309]. Case [167, 229, 317, 236, 177, 178, 287, 316, 187, 315, 105, 319, 291, 293, 311, 150, 214, 297, 157, 222, 273, 275, 312, 313, 62, 280, 281, 168, 131, 92, 133, 180, 244, 250, 78, 96, 188, 198, 199, 324, 254, 146, 202, 203, 108, 205, 206, 148, 208, 152, 158, 217, 88, 303, 123, 308, 226, 165, 321, 166, 279]. Case-Based [229, 291, 293, 311, 167, 317, 177, 178, 316, 315, 319, 150, 214, 297, 157, 222, 275, 62, 280, 281, 168, 131, 92, 133, 180, 244, 78, 96, 188, 199, 324, 254, 146, 202, 203, 108, 205, 206, 148, 152, 158, 217, 88, 123, 165, 321, 166, 279]. case-memory [250]. cases [232, 173, 174, 235, 138, 259, 260, 151, 296, 306]. CASSIOPÉE [282]. Cataloging [194]. Categorization [21, 196]. categorized [214]. causal [56]. CBR [170, 284, 288]. Center [309]. century [57]. CFM [282]. Chambery [320]. change [235]. Chantilly [324]. characteristics [281]. characters [110]. CHARADE [219]. CheckL [257]. CheckUp [257]. Chemical [87, 175, 183]. Choice [20]. city [301]. clarification [66]. Classification [54, 124]. Clique [64, 63]. clustering [286]. Cognition [107]. cognitive [38, 320]. color [154]. Coloured [262]. columns [85]. common [172, 73]. CommonKADS [91, 182]. Communication [132]. Comparing [185]. Comparison [239, 175, 193]. Competence [274]. competent [126]. Compilation [91]. complex [247, 150, 119, 306]. complexity [38]. Components [182]. composite [266]. Computational [76, 215]. Computer [31, 309, 241, 243, 145]. computer-aided [243]. Computer-Support [309]. Computerunterstützung [106, 144, 323]. Computing [197, 212]. Concept [54, 221]. Concepts [262, 189, 98, 99, 205]. Conceptual [238, 286, 78, 112, 266, 118]. concrete [85]. Conf [314, 100]. Conference [322, 273, 320]. confluence [134]. constraint [149, 75]. constraint-based [75]. construction [283, 145, 118, 303]. content [192, 154]. continuous [26]. Control [55, 259, 260]. convictions [285]. cooperation [102]. Cooperative [132, 303, 162]. copy [37]. correct [252]. Cranfield [8]. critics [162].

d’Architecture [2]. data [172, 194]. Databases [87, 183, 136, 216]. Datenbanken [53]. Datensammlungen [248]. day [38]. DC [317, 316, 319]. Decision [280]. DEJAVU [131]. densities [110]. descriptions [98]. Design [5, 20, 132, 240, 55, 8, 100, 322, 105, 293, 311, 72, 76, 169, 231, 232, 283, 131, 172, 173, 174, 285, 24, 176, 133, 177, 178, 237, 241, 242, 184, 187, 78, 96, 188, 52, 195, 145, 201, 255, 146, 256, 257, 259, 260, 261, 148, 150, 266, 215, 297, 158, 217, 302, 123, 275, 276, 277, 312, 75, 234]. Design-Domänen [234]. designer [174]. designs [247, 198, 320]. Desk [309]. destructive [157]. detailing [85]. Detecting [220, 25]. Determining [8]. deutschsprachigen [310]. Development [95, 161, 273, 39, 114]. Diagnosesituationen [89]. diagnosis [181, 200, 56]. Diagnostik [292, 218]. Diagnostik-Shell [218]. diagrams [198]. different [189]. Digital [15, 194, 145]. Digitales [323, 106, 144]. Dimensional [87, 175, 183]. directional [110]. Directions [291]. disciplinary [184]. disease [17]. distance [98]. distances [197, 212]. Distributed [264, 149, 213]. documents [277]. DOM [233, 231, 232, 283, 296]. DOM-ARCADE [283]. domain [281, 233, 201, 279]. domains [168, 173]. Domäne [207]. Domänen [234, 77]. données [2]. dot [124]. driven [164]. Dynamic [208, 210]. Dynamik [267]. dynamischen [267].

early [124]. Ecole [2]. Edifices [1]. edition [71]. effective [192]. Efficient [192, 59, 286]. Effiziente [163]. einem [135]. Einführung [48]. elicitation [275]. emergencies [219]. Empirie [111]. empirische [310]. empirischen [34]. Energy [59]. Engineering [290, 264, 162, 318, 231, 243, 303, 123, 130, 230]. Entscheidungsunterstützung [292]. Entscheidungsunterstützungssysteme [310]. Entwerfen [298]. Entwicklung [207, 33, 53, 227]. Entwicklungssystem [295]. Entwurf [106, 144, 323, 272]. Environment [95, 232]. episodes [244]. Episodisches [74]. Erfahrungswissen [43]. Erkennungssystem [160]. Erste [50]. Erstellung [227]. European [324, 321]. Evaluating [206]. evaluation [283, 304, 289]. event [143]. EVENTS [143]. EVENTS-II [143]. EWCBR [324, 321]. EWCBR-94 [324]. Exact [252]. examples [116]. exekutiver [33]. Exemplar [54]. Exemplar-Based [54]. Experiences [291, 285]. Experiment [50, 23]. experimentelle [48]. experise [38]. Expert [95, 266, 45, 222, 162]. Experte [94]. Expertensysteme [83]. Expertensystemen [90, 43]. Expertensystemkern [77]. Expertensystemshells [121]. Expertise [182, 193, 82, 103, 60, 51, 84]. experts [21, 303]. explanation [116]. Explicit [115]. Exploration [76]. Exploring [146]. Extensions [7, 287]. extensive [145].

FABEL

[285, 201, 269, 271, 224, 225, 305, 312]. FAC [81]. Factors [8]. FAENSY [295]. fall [295]. fall- [295]. Fallbasierte [93, 292, 121]. fallbasierten [101, 207, 268, 272]. fallbasierter [267, 227]. fallbasiertes [74, 309, 310, 90, 289, 129]. Fallbasis [270]. Fallbibliotheken [97]. Fälle [268]. Fällen [101, 163]. fallfokussierenden [89]. Fallgestütztes [298]. fallorientiertem [43]. feature [110]. features [122, 29]. Fehlerdiagnose [282]. FERN [116]. fertiges [142]. fibre [266]. Figural [9]. Figural-Wahrnehmung [9]. files [175]. Finding [138]. First [273, 19, 222, 321, 5]. first-order [19]. Fischen [300]. flat [212]. Florida [315]. Flugzeuge [282]. Form [10, 122]. Formal [91, 189, 139, 82, 103]. formalization [19]. Formalizations [108, 180]. formalize [193]. Formalizing [217]. Forschung [34]. foundation [208]. foundational [167]. fragment [175]. fragment-based [175]. fragments [185]. framework [281, 313]. France [324, 320]. freely [110]. Full [66]. Full-text [66]. fundierten [67]. Further [17, 66]. Future [291]. Fuzzy [186, 203].

ganzheitliche [86]. gap [169]. Gebäude [265]. Gebäudehülle [58]. Gebäuden [106, 144, 323]. Geburtstag [13]. geheimnisvollere [30]. General [249]. generation [222]. generic [296]. Generierung [253]. genre [1]. geographic [194]. Geometric [199, 249]. Germany [321]. Gestalt [3, 210, 4, 160]. Gestalt-Erkennungssystem [160]. Gestalts [221, 220]. Gestalttheorie [13]. Giving [139]. Global [57]. Glossar [41]. GO [207]. GO-Domäne [207]. Grafikdesigner [120]. Grammars [31]. Graph [171, 31]. Graph-based [171]. großer [97]. groundwork [196]. grouping [208, 210, 29]. Grundlagen [69, 48, 129]. guidance [133, 179, 210]. guidelines [313].

Handlungskontrolle [30]. handwritten [110]. Haustechnik [42]. having [172]. held [317, 316, 322]. Help [309]. Help-Desk- [309]. hermeneutisch [67]. Heuristic [149, 299]. hierarchical [244, 26]. high [119]. histories [166]. hochautomatisierter [265]. Holiday [315]. home [200]. Hotel [317, 316]. house [52]. Human [60]. Hybrid [202, 170, 224]. hypermedia [258].

ICCBR [273]. ICCBR-95 [273]. idea [68, 271]. Ideas [20, 287]. II [143]. im [245, 288, 153, 155, 295, 159, 53, 268, 310]. Image [137, 192, 157, 216]. images [185, 211, 113, 154]. imagination [174]. implement [213]. Implementierung [272]. important [172]. Improving [165]. INCOME [117]. Indexing [8, 170, 246, 247, 211, 152]. indices [220]. Induction [151]. Industrial [229, 240]. inference [109]. Inferenzprozessen [190]. information [24, 66, 186, 251, 194, 23, 304]. Informationsstruktur [53]. Informationssysteme [117]. Informationsverarbeitung [46]. informed [286]. infrastructure [194]. inkrementelles [230]. Inn [315]. Innovation [176, 76, 204]. innovative [234]. Inquiry [290]. installationsmodell [35]. Integralen [42]. integrate [189]. integrated [213, 302]. Integration [289, 295]. integrierten [53]. integrierter [12]. Integritätsbedingungen [140]. Intelligence [100, 322, 320]. Intelligent [32, 59, 276, 256]. Intelligente [309]. intensive [62, 208, 166]. Interaction [296]. Interactionist [107]. Interactive [241, 195]. interaktiver [117]. interdisciplinary [320]. Internat [100]. International [31, 322, 273, 320]. Interpretation [184, 92, 157, 301]. interpreting [80]. Introduction [37, 16, 209, 156, 158, 27]. invariant [122]. Invention [204]. Investigation [87]. IQ [38]. isomorphism [17]. Issues [311, 158, 167, 211].

Japanese [110]. Joint [320]. judging [25]. July [319]. June [315].

KADS [127, 193, 102, 82, 103, 161, 130]. Kaiserslautern [321]. KARL [79, 139]. KI [234]. KI-Methoden [234]. KICS [166]. kit [188, 118]. Klassenformat [245]. Klassifizieren [289]. Kleines [41]. Knowledge [62, 91, 54, 55, 104, 290, 115, 161, 172, 178, 242, 79, 52, 259, 147, 208, 19, 113, 114, 85, 296, 215, 303, 164, 130, 166, 75, 230]. Knowledge-Based [161, 178, 52, 114, 85, 164]. Knowledge-intensive [62, 208, 166]. kognitiver [33]. Komplexe [155]. konfigurierbare [218]. Konfigurierungsaufgaben [77]. Konstrukte [50]. Konstruktion [106, 144, 323]. Kontextsensitivität [111]. Kontextualisierung [84]. konzeptuelle [125].

labels [68]. Lageplänen [253]. Language [73, 235, 79, 82, 103]. languages [193, 114]. LanPlan [258]. large [138, 216]. Lausanne [322]. law [152]. layout [237, 70, 248]. layouts [283, 246, 269, 223, 224, 225]. learned [98]. Learning [54, 281, 177, 178, 6]. Lecons [2]. legal [299]. Lehrbuch [34]. Lehrer [94]. Leistungen [33]. Lernen [89]. Lessons [291]. level [104, 147, 119]. Library [182]. Life [49, 57]. line [26, 110]. Lisp [73]. literal [18]. literature [194]. logic [203]. Lösung [64].

MAC [81]. MAC/FAC [81]. Machine [5, 72]. Madison [317, 316]. make [284]. making [127]. Man [7, 270]. Management [86, 145, 219]. manufacturing [254]. many [68]. map [301]. maschinelles [160]. Maschinenbegriffs [67]. matching [98].

MAX [160]. maximum [63, 64]. Maximum-Clique-Problems [64]. May [317, 316, 315]. measures [175]. mechanical [131]. Mechanismen [30, 159]. Media [7, 57]. Medical [56]. medieninstallation [12]. mehrgeschossige [12]. memory [250, 197, 212, 27]. Mengen [141]. Metakognition [30]. Metaphor [107]. Metaphorische [204]. method [122]. Methoden [289, 267, 121, 234]. methodischen [48]. methodological [167]. Methodology [91, 242, 23]. Methods [91, 262, 115, 213, 87, 281, 168, 68, 226]. metrics [199]. Metzger [13]. MIDI [12]. MIKE [230]. MIKE-Ansatz [230]. mindful [278]. mittels [153]. Model [181, 164, 76, 92, 176, 81, 255, 116, 118, 215, 147, 127]. Model-based [181, 176]. Model-driven [164]. MODEL-K [147, 127]. Modeling [210, 99, 143, 114, 164, 75]. Modell [106, 144, 323, 268, 50]. Modell-Konstrukte [50]. Modellbasiertes [230]. Modelle [83, 125]. Modellierung [50, 111]. Modelling [91, 182, 130]. Models [68, 51, 193, 82, 103]. modernen [13]. Modernes [1]. modules [239]. MoMo [290, 128]. Monitoring [190]. MOPs [27]. morgen [58]. most [172]. Multi [45, 184, 19]. multi-agent [19]. multi-disciplinary [184]. Multi-expert [45]. Multidimensional [250]. multiple [258, 152, 226]. Muratoris [41]. Mustererkennung [159].

natural [6]. need [174]. Nets [262, 287, 119]. network [56]. Netze [69, 65]. Netzen [140]. Neuronale [69]. noch [30]. noisy [98]. non [157]. non-destructive [157]. Nonstandard [205]. November [324, 321]. novices [21]. nützlicher [268]. Nutzung [97].

object [98, 119, 75]. object-oriented [75]. objects [228, 11, 25, 184, 29]. Objekte [153]. objektorientierten [135]. offenes [12]. ohne [191, 263, 270]. ökologischen [267]. On-line [110]. one [68]. ontology [233]. Open [95, 168, 226, 105]. operational [139, 114]. operationalize [193]. Operators [261]. order [19]. organisation [215]. organization [27]. Oriented [290, 173, 242, 75]. Osnabrück [31].

P [88]. panels [266]. Papers [100, 324, 319, 321]. Parallèle [1]. Part [24]. partially [98]. Patdex [89]. Patdex/2 [89]. Pattern [16, 71, 37, 122]. pattern-specific [122]. patterns [124]. Pensacola [315]. Perception [10, 25, 28]. Perceptual [29, 124]. Performance [8]. perils [80]. perspectives [304]. Petri [65, 140, 262, 119]. Petri-Netze [65]. Petri-Netzen [140]. physics [21]. Phytoplankton [267]. Phytoplankton-Dynamik [267]. Pictorial [251]. Picture [136, 15, 134]. Picture-queries [136]. pipes [252]. placement [252]. PLAKON [77]. PLAKON-Buch [77]. planning [235, 206, 19, 301]. plans [220, 221]. Plants [59]. Planung [42]. Planungs [77]. Planungs- [77]. platform [219]. PM5 [255]. points [138]. Polytechnique [2]. Portable [95]. Practical [262]. practice [284]. Praktikum [48]. praktischen [48]. precedent [215]. Precedents [215]. Precis [2]. precision [80]. prediction [181]. prerequisite [172]. Principled [161]. Principles [4, 210]. probabilistic [56]. Problem [182, 55, 40, 105, 115, 28, 281, 63, 244, 149]. Problem-Solving [115, 40]. Problemlösen [46, 191]. problems [21, 109, 152, 275, 306, 64]. Proc [314]. Proceedings [322, 317, 316, 315, 320]. Produkt [142]. professionellen [94]. Profiles [271]. Prognose [272, 267].

Programmentwurf [117]. Programmierstil [245]. Programmierwerkstatt [67]. project [285, 154, 8]. PROSET [141]. PROSET-Ansatz [141]. Prototyp [142]. prototype [164]. prototype-driven [164]. Prototyping [314, 135, 318, 147, 164, 141]. providing [196]. Prozeß [267]. Prozesse [33]. Psychologie [94, 13, 48]. Psychology [4, 3]. Psychometrics [14].

QBIC [154]. Qualitätsstrategie [86]. Quality [232, 86, 259, 260]. queries [136]. querying [192, 154].

races [38]. Raum [310]. räumliche [155]. real [11]. realizations [257]. reasoner [203]. Reasoning [229, 291, 293, 311, 273, 61, 62, 167, 131, 317, 92, 180, 286, 244, 316, 199, 315, 324, 254, 202, 108, 205, 147, 319, 113, 150, 151, 214, 152, 157, 88, 275, 165, 321, 166, 279]. recall [80]. reciprocal [134]. Recognition [16, 26, 37, 137, 143, 110, 122]. Recueil [1]. reenforced [85]. refining [235]. reflection [104]. reflective [126]. regelbasierten [295]. Regelung [265]. regulations [166, 279]. reinforced [266]. reinterpretation [208]. Relating [281]. relational [25]. relationale [53]. Relationen [155]. ReMind [170]. Reminding [27]. Repeat [294]. Representation [54, 235, 21, 79, 148, 113, 75]. representations [197, 258, 212]. Representing [173]. Requirements [318, 169]. Research [8, 273]. Results [24]. Retrieval [44, 22, 269, 223, 224, 225, 277, 170, 24, 66, 134, 287, 137, 246, 247, 249, 186, 251, 81, 203, 32, 113, 299, 23, 304, 275, 226, 165, 263]. Reusable [182]. Reuse [290, 320, 246, 247, 277]. Reuse-Oriented [290]. reusing [244, 313]. Review [229, 194]. revisited [200]. role [196, 114]. roles [68]. Rolle [288]. rotation [122]. Roude [261]. routine [261]. run [127].

satisfaction [149]. scale [122]. Scenario [302]. Scene [25]. scenes [11]. scheduler [213]. schemata [28]. Schließen [90, 272, 129, 309, 310, 268, 74]. Schließens [101]. Schritte [50]. Science [31]. search [149, 29]. Searching [11, 183, 87, 175, 216]. Second [324, 72, 222]. Sehsystem [159]. Selbststeuerung [30]. selected [324, 321]. semantics [139]. semantische [204]. services [283]. Shaker [49]. shape [175, 154]. sharing [172]. Shell [218]. shells [222]. shift [122]. similar [134, 269, 223, 224, 225]. similar-picture [134]. Similarity [175, 14, 148, 87, 216, 88, 61, 171, 133, 179, 189, 39, 81, 196, 199, 47, 205, 208, 113, 18, 221]. similarity-based [81]. simple [92]. simulation [6]. Software [264, 67]. Software-Werkzeuge [67]. Softwarewerkzeuge [83]. solve [306]. Solving [182, 55, 105, 115, 281, 244, 40, 28]. Some [211, 304]. Spatial [113]. special [37]. specialists [312]. specific [122]. specification [119]. Stadtmorphologie [41]. Standgewässern [272]. statutory [166]. step [203, 165]. stereo [200]. Steuerung [265]. strategic [147]. Strategies [55, 59, 40]. strategy [170]. Stratified [244]. string [113]. Structural [133, 307, 171, 179, 98, 47]. structure [109, 112, 124]. structured [139]. Structures [55, 87, 175, 183, 26]. structuring [250]. strukturiert [270]. Strukturierte [97]. Study [105, 24, 38, 308]. Suche [268, 263]. Suchen [191]. Supervised [124]. Support [309, 280, 174, 285, 178, 240, 145, 301, 302]. supported [233]. supporting [299]. Surface [47, 39]. survey [313]. sustained [62]. Switzerland [322]. Syn* [237]. Syntactic [16, 37, 98]. synthesis [180, 123].

System

[243, 190, 161, 167, 231, 285, 242, 52, 102, 12, 26, 22, 206, 19, 85, 217, 222, 166, 75, 89]. system-user [102]. Systematicity [39]. Systeme [309, 50, 227]. Systemen [101]. Systems [95, 8, 207, 272, 170, 284, 45, 114, 158, 219, 162, 304, 126, 164, 308, 313].

task [281, 242]. task-oriented [242]. tasks [257, 68, 150]. Tcl [156]. technical [277]. Techniken [83]. techniques [44, 183, 139, 301]. technischen [77, 89]. technology [169, 194]. Term [179]. Term-based [179]. test [168]. testing [157]. tests [22]. text [66]. texture [154]. their [185, 31]. Theory [5, 72, 24]. Third [322]. Thirteenth [320]. Three [87, 175, 183]. Three-Dimensional [87, 175, 183]. Tk [156]. tool [188, 123]. Tools [229, 271]. Topics [321]. TOPO [249, 307]. topologic [249]. topologies [246, 247]. Total [86]. tout [1]. transfer [47]. transformations [57]. tree [134]. Triebwerken [282]. tutor [258].

Überblick [129]. Ubungen [48]. Umsetzung [50]. uncertainty [88]. undergrouping [25]. Understanding [7, 172, 71]. Unified [54]. unit [134]. Untersuchung [310]. Use [262, 279]. used [261, 213, 220]. user [102]. Using [198, 259, 260, 173, 236, 258, 110, 113, 56, 154, 269, 301, 122].

value [197, 212]. values [80]. variations [167]. Verfahren [64, 295]. Versenken [300]. very [224]. views [112]. Village [57]. violations [25]. Visual [10, 174, 29]. visualisation [173]. visualisation-oriented [173]. VLSI [70]. vol [262]. Volume [8]. Vorauswahl [163].

Wahrnehmung [9].

Wahrnehmungspsychologie [120].

Washington [317, 316, 319]. Wechselnde [270]. Weg [142]. weitgehende [106, 144, 323]. Werkzeug [67]. Werkzeug- [67]. Werkzeuge [67, 121]. which [174, 193]. Wie [270]. Wiederverwendbare [218]. Wissen [74, 270]. Wissens [94]. Wissensakquisition [93, 83]. Wissensbasierte [265]. wissensbasierter [50]. Wissensbasis [50]. within [232]. without [137]. Wolfgang [13]. wonach [270]. Work [49, 284]. Workbench [95]. Working [314]. Workshop [316, 31, 315, 324, 320, 317, 319, 321]. world [168, 11, 57]. Worlds [105].

Ziel [142]. Zugriffsstrukturen [163]. zum [248, 190, 13, 272, 89]. zur [64, 43, 94, 268, 50, 227, 295].

References

Durand:1801:RPE

[1] J. N. L. Durand. Recueil et Parallèle des Edifices de tout genre Anciens et Modernes. ????, ????, 1801. Paris An IX, dit “le grand Durand”.

Durand:1802:PLA

[2] J. N. L. Durand. Precis des Leçons d’Architecture données à l’Ecole Polytechnique. ????, ????, 1802. Paris An X.

Kohler:1929:GP

[3] W. Köhler. Gestalt Psychology. Liveright, New York, 1929.

Kofka:1935:PGP

[4] K. Kofka. Principles of Gestalt Psychology. Harcourt Brace, New York, 1935.

Banham:1960:TDF

[5] Peter Reyner Banham. Theory and Design in the First Machine Age. ????, ????, 1960. 338 pp. Reprinted 1980 by MIT Press, Cambridge, MA.

Feigenbaum:1963:SNL

[6] Edward A. Feigenbaum. The simulation of natural learning behaviour. In E. A. Feigenbaum and J. Feldman, editors, Computers and Thought, pages 297–309. McGraw-Hill, New York, NY, USA, 1963.

McLuhan:1964:UME

[7] Marshall McLuhan. Understanding Media: the Extensions of Man. McGraw-Hill, New York, NY, USA, 1964. ISBN 0-262-63159-8. xxiii + 365 pp. LCCN P90 .M26 1994. Reprinted by MIT Press, Cambridge, MA, 1994.

Cleverdon:1966:FDP

[8] Cyril Cleverdon, Jack Mills, and Michael Keen. Factors Determining the Performance of Indexing Systems: ASLIB Cranfield Research Project. Volume 1: Design. ASLIB Cranfield Research Project, Cranfield, 1966. 377 pp.

Metzger:1966:FW

[9] W. Metzger. Figural-Wahrnehmung. In W. Metzger, editor, Handbuch der Psychologie, Bd. 1 Allgemeine Psychologie, 1.1. Der Aufbau des Erkennens, pages 693–744. Hogrefe, Göttingen, 1966.

Zusne:1970:VPF

[10] L. Zusne. Visual Perception of Form. Academic Press, New York, USA, 1970.

Biederman:1973:SOR

[11] I. Biederman, A. L. Glass, and E. W. Stacy. Searching for objects in real world scenes. Journal of Experimental Psychology: General, 91:22–27, 1973. CO-DEN JPGEDD. ISSN 0096-3445 (print), 1939-2222 (electronic).

Haller:1974:MOS

[12] Fritz Haller. MIDI — ein offenes system für mehrgeschossige bauten mit integrierter medieninstallation. USM bausysteme haller, Münsingen, 1974. ISBN ???? LCCN ????

Ertel:1975:GMP

[13] Suitbert Ertel, Lilly Kemmler, and Michael Stadler, editors. Gestalttheorie in der modernen Psychologie: Wolfgang Metzger zum 75. Geburtstag. Dr. Dietrich Steinkopff Verlag, Darmstadt, Germany, 1975. ISBN 3-7985-0400-8. ix + 319 pp. LCCN BF203.G47.

Gregson:1975:PS

[14] Robert A. M. Gregson. Psychometrics of Similarity. Academic Press, New York, USA, 1975. ISBN 0-12-301550-2. ix + 262 pp. LCCN BF39 .G73.

Rosenfeld:1976:DPA

[15] Azriel Rosenfeld. Digital Picture Analysis, volume 11 of Topics in Applied Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07579-8. xi + 351 pp. LCCN TA1630 .D53.

Gonzalez:1978:SPR

[16] Rafael C. Gonzalez and Michael G. Thomason. Syntactic Pattern Recognition: an Introduction. Addison-Wesley, Reading, MA, USA, 1978. ISBN 0-201-02931-6. 283 pp.

Gati:1979:FAB

[17] Georg Gati. Further annotated bibliography on the isomorphism disease. J. Graph Theory, 3:95–109, 1979. CODEN JGTHDO. ISSN 0364-9024.

Ortony:1979:BLS

[18] Andrew Ortony. Beyond literal similarity. Psychological Review, 86:161–180, 1979.

Konolige:1980:FOF

[19] Kurt Konolige. A first-order formalization on knowledge and action for a multi-agent planning system. Technical Report 232, SRI International, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, Tel: +1 415 859 6387, FAX: +1 415 859-6028, 1980.

Banham:1981:DCI

[20] Peter Reyner Banham and Penny Sparke. Design by Choice: Ideas in Architecture. Academy Editions, London, 1981. ISBN ???? 152 pp. LCCN NA680.B316 1981.

Chi:1981:CRP

[21] M. T. H. Chi, P. Feltovich, and R. Glaser. Categorization and representation of physics problems by experts and novices. Cognitive Science, 5:121– 152, 1981.

Jones:1981:RST

[22] Karen Sparck Jones. Retrieval system tests 1958 — 1978. In Karen Sparck Jones, editor, Information Retrieval Experiment, pages 213–255. Butterworths, London, UK, 1981. ISBN 0-408-10648-4.

Robertson:1981:MIR

[23] Stephen E. Robertson. The methodology of information retrieval experiment. In Karen Sparck Jones, editor, Information Retrieval Experiment, pages 9–31. Butterworths, London, UK, 1981. ISBN 0-408-10648-4.

Belkin:1982:AIR

[24] Nicklas J. Belkin, Robert Oddy, and Helen M. Brooks. ASK for information retrieval. Part 1: Background and theory. Part 2: Results of a design study. Journal of Documentation, 38: 61–71, 145–164, 1982. CODEN JDO-CAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Biederman:1982:SPD

[25] I. Biederman, R. J. Mezzanotte, and J. C. Rabinowitz. Scene perception: detecting and judging objects undergrouping relational violations. Cognitive Psychology, 14:143–177, 1982.

Hartmann:1982:RCL

[26] Georg Hartmann. Recognition of continuous line structures by a hierarchical system. In Proc. 6th Internat. Conf. on Pattern Recognition (ICPR), pages 195– 200. IEEE Computer Society, New York, München, October 1982.

Schank:1982:RMO

[27] Roger C. Schank. Reminding and memory organization: an introduction to MOPs. In Wendy G. Lehnert and Martin H. Ringle, editors, Strategies for Natural Language Processing, pages 455– 493. Erlbaum, Hillsdale, NJ, 1982. ISBN 0-89859-266-6.

Silver:1982:PPP

[28] E. A. Silver. Problem perception, problem schemata and problem solving. Journal of Mathematical Behavior, Sum 3(2):169–181, 1982. ISSN 0732-3123.

Treisman:1982:PGA

[29] A. Treisman. Perceptual grouping and attention in visual search for features and for objects. Journal of Experimental Psychology: Human Perception and Performance, 8(2):194–214, 1982.

Brown:1983:MHS

[30] A. L. Brown. Metakognition, Handlungskontrolle, Selbststeuerung und andere noch geheimnisvollere Mechanismen. In F. E. Weinert and R. H. Kluwe, editors, Metakognition, Motivation und Lernen, pages 60–80. Kohlhammer, Stuttgart, 1983.

Ehrig:1983:GGT

[31] Hartmut Ehrig, Manfred Nagl, and Grzegorz Rozenberg, editors. Graph Grammars and their Application to Computer Science: 2nd International Workshop, Osnabrück, 1982. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 3-540-12310-5. 452 pp.

Jones:1983:IR

[32] Karen Sparck Jones. Intelligent retrieval. In Kevin P. Jones, editor, Intelligent Information Retrieval: Informatics 7, pages 136–142. Aslib, London, Cambridge, March 1983.

Kluwe:1983:EEP

[33] R. H. Kluwe and K. Schiebler. Entwicklung exekutiver Prozesse und kognitiver Leistungen. In F. E. Weinert and R. H. Kluwe, editors, Metakognition, Motivation und Lernen, pages 31–59. Kohlhammer, Stuttgart, 1983.

Bortz:1984:LEF

[34] J. Bortz. Lehrbuch der empirischen Forschung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-12852-2. 649 pp.

Haller:1985:AI

[35] Fritz Haller. ARMILLA — ein installationsmodell. Institut für Baugestaltung, Universität Karlsruhe, Karlsruhe, Germany, 1985. ISBN ???? 141 pp. LCCN ????

Vincent:1985:AB

[36] Gordon S. Vincent and John Peacock. The Automated Building. Architectural Press, London, 1985. ISBN 0-85139-737-9. 211 pp. LCCN TH6012.V56 1985. Additional material by Alec Robertson.

Bunke:1986:ISC

[37] Horst Bunke and Alberto Sanfeliu. Introduction [to a special copy on syntactic pattern recognition]. Pattern Recognition, 19:249–253, 1986. CODEN PT-NRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Ceci:1986:DRS

[38] S. J. Ceci and J. K. Liker. A day at the races: a study of IQ, experise and cognitive complexity. Journal of Experimental Psychology: General, 115:255–266, 1986. CODEN JPGEDD. ISSN 0096-3445 (print), 1939-2222 (electronic).

Gentner:1986:SSS

[39] Dedre Gentner and C. Toupin. Systematicity and surface similarity in the development of analogy. Cognitive Science, 10:277–300, 1986.

Gick:1986:PSS

[40] Mary L. Gick. Problem-solving strategies. Educational Psychologist, 21(1 & 2):99–120, 1986.

Malfroy:1986:KGM

[41] S. Malfroy. Kleines Glossar zu Muratoris Stadtmorphologie. ARCH+, Zeitschrift für Architektur und Städtebau, ??:66–73, June 1986.

Suter:1986:HIP

[42] P. Suter, R. Geller, N. Kohler, and J. van Gilst. Haustechnik in der Integralen Planung. Impulsprogramm Haustechnik, Bundesamt für Konjunkturfragen, Bern, Schweiz, 1986.

Bartsch-Sporl:1987:ABF

[43] Brigitte Bartsch-Spörl. Ansätze zur Behandlung von fallorientiertem Erfahrungswissen in Expertensystemen. KI — Künstliche Intelligenz, 1(4):32–36, 1987.

Belkin:1987:RT

[44] Nicholas J. Belkin and W. Bruce Croft. Retrieval techniques. In Martha E. Williams, editor, Annual Review of Information Science and Technology, volume 22, pages 109–145. Elsevier, Amsterdam, The Netherlands, 1987. ISBN 0-444-70302-0.

David:1987:MES

[45] B. T. David. Multi-expert systems for CAD. In tenHagen et al., editors, Intelligent CAD Systems. Springer, Berlin, Noordwijkerhout, 1987.

Dorner:1987:PI

[46] Dietrich Dörner, editor. Problemlösen als Informationsverarbeitung. Kohlhammer, Stuttgart, 1987. ISBN 3-17-001353-X. 151 pp. LCCN BF441.D63.

Holyoak:1987:SSS

[47] K. J. Holyoak and K. Koh. Surface and structural similarity in analogical transfer. Memory and Cognition, 115:332– 340, 1987. CODEN MYCGAO. ISSN 0090-502X.

Luer:1987:AEP

[48] Gerd Lüer. Allgemeine experimentelle Psychologie: eine Einführung in die methodischen Grundlagen mit praktischen Ubungen für das experimentelle Praktikum. Fischer, Stuttgart, 1987. ISBN 3-437-00477-8. xiii + 478 pp. LCCN ????

Sprigg:1987:SLW

[49] June Sprigg. Shaker — Life, Work, and Art. Stewart, Tabori & Chang, New York, 1987. ISBN 1-55670-011-3. 272 pp. LCCN BX9771.S68 1987.

Voss:1987:EUB

[50] Angi Voß, Bernd S. Müller, and Jürgen Walther. Über ein Experiment zur Umsetzung einer BABYLON-Wissensbasis in Modell-Konstrukte: Erste Schritte zur Modellierung wissensbasierter Systeme. In W. Brauer and W. Wahlster, editors, Wissensbasierte Systeme: 2. Internationaler GI-Kongreß, pages 358– 368. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 3-540-18494-5.

Wielinga:1987:ME

[51] Bob J. Wielinga and Joost A. Breuker. Models of expertise. In Ben du Boulay, David Hogg, and Luc Steels, editors, Advances in Artificial Intelligence — II, ECAI-86, Brighton, pages 306–318. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987.

Garijo:1988:KBS

[52] F. J. Garijo and L. A. de Garrido. A knowledge-based system for house design. In Proceedings of the 1988 IEEE International Conference on Systems, Man, and Cybernetics, pages 806–809. IEEE Computer Society Press, International Academic Publishers, Beijing, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1988. ISBN 7-80003-039-3.

Richter:1988:EII

[53] P. Richter. Entwicklung einer integrierten Informationsstruktur für relationale Datenbanken im Bauwesen. PhD thesis, Gesamthochschule Kassel, Kassel, Germany, 1988.

Bareiss:1989:EBK

[54] R. Bareiss. Exemplar-Based Knowledge Acquisition: a Unified Approach to Concept Representation, Classification and Learning. Academic Press, New York, USA, 1989. ISBN 0-12-078268-X (??invalid ISBN checksum??). 169 pp.

Brown:1989:DPS

[55] David C. Brown and B. Chandrasekaran. Design Problem Solving: Knowledge Structures and Control Strategies. Pitman Publishing Ltd., London, UK, 1989. ISBN 0-273-08766-5. 199 pp.

Long:1989:MDU

[56] William J. Long. Medical diagnosis using a probabilistic causal network. Applied Artificial Intelligence, 3:367–383, 1989.

McLuhan:1989:GV

[57] Marshall McLuhan and Bruce R. Powers. The Global Village: transformations in world life and media in the 21st century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-507910-8. xiii + 220 pp. LCCN ????

Muller:1989:GM

[58] H. Muller. Die Gebäudehülle von morgen. In J. Gauchel, editor, KI-Forschung im Baubereich. Universität Karlsruhe, 1989.

Payne:1989:SEE

[59] F. William Payne. Strategies for Energy Efficient Plants & Intelligent Buildings. Fairmont Press, Lilburn, GA, 1989. ISBN 0-88173-020-3, 0-13-850686-8 (Prentice-Hall). xv + 603 pp. LCCN TJ163.5.P68 S77 1987. Based on papers presented at the 9th World Energy Engineering Congress, Oct. 21–24, 1986 in Atlanta, GA.

Reimann:1989:HE

[60] P. Reimann and M. T. H. Chi. Human expertise. In K. J. Gilhooly, editor, Human and Machine Problem Solving, pages 161–191. Plenum Press, New York, NY, USA; London, UK, 1989.

Vosniadou:1989:SAR

[61] Stella Vosniadou and Andrew Ortony. Similarity and Analogical Reasoning.

Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-38935-6. 592 pp.

Aamodt:1990:KIC

[62] Agnar Aamodt. Knowledge-intensive case-based reasoning and sustained reasoning. In L. C. Aiello, editor, Proceedings of ECAI-90, pages 1–6. Pitman, London, Stockholm, 1990. ISBN 0-273-08822-X.

Babel:1990:BBA

[63] Luitpold Babel and Gottfried Tinhofer. A branch and bound algorithm for the maximum clique problem. ZOR — Mathematical Methods of Operations Research, 34:207–217, 1990. CODEN ZMRREP. ISSN 0340-9422.

Babel:1990:BBV

[64] Luitpold Babel. Ein Branch and Bound-Verfahren zur Lösung des Maximum-Clique-Problems. Diss., Technische Universität München, München, Germany, December 1990. 114 pp.

Baumgarten:1990:PN

[65] Bernd Baumgarten. Petri-Netze. BI-Wissenschaftsverlag, Mannheim, Germany; Wien, Austria; Zürich, Switzerland, 1990. ISBN 3-411-14291-X. 369 pp.

Blair:1990:FTI

[66] David C. Blair and M. E. Maron. Fulltext information retrieval: further analysis and clarification. Information Processing and Management, 26:437–447, 1990. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Budde:1990:SWP

[67] Reinhard Budde and Heinz Züllighoven. Software-Werkzeuge in einer Programmierwerkstatt: Ansätze eines hermeneutisch fundierten Werkzeugund Maschinenbegriffs. Number 182 in GMD-Bericht. R. Oldenbourg, München, Germany, 1990. ISBN 3-486-21629-5. 288 pp. Also: Berlin, Techn. Univ., Diss., 1989.

Karbach:1990:MMR

[68] Werner Karbach, Marc Linster, and Angi Voß. Models, methods, roles and tasks: many labels — one idea? Knowledge Acquisition, 2:279–299, 1990.

Kratzer:1990:NNG

[69] Klaus Peter Kratzer. Neuronale Netze: Grundlagen und Anwendungen. Carl Hanser Verlag, München; Wien, 1990. ISBN 3-466-16026-4 (??invalid ISBN checksum??). 211 pp.

Kuh:1990:RAV

[70] Ernest S. Kuh and Tatsuo Ohtsuki. Recent advances in VLSI layout. Proceedings of the IEEE, 78:237–263, 1990. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Niemann:1990:PAU

[71] H. Niemann. Pattern Analysis and Understanding, 2nd edition. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2 edition, 1990. ISBN 3-540-51378-7. 371 pp.

Pawley:1990:TDS

[72] Martin Pawley. Theory and Design in the Second Machine Age. Blackwell, Oxford, 1990. ISBN 0-631-15828-6. xii + 189 pp. LCCN NA2500 .P385 1990.

Steele:1990:CLL

[73] Guy L. Steele Jr. Common Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1990. ISBN 1-55558-041-6 (paperback), 1-55558-042-4 (hardcover), 0-08-050226-1 (e-book), 0-13-152414-3 (Prentice-Hall). xxiii + 1029 pp. LCCN QA76.73.L23 S73 1990. With contributions by Scott E. Fahlman and others, and with contributions to the second edition by Daniel G. Bobrow and others. See also [?, ?].

Strube:1990:EWF

[74] Gerhard Strube and Dietmar Janetzko. Episodisches Wissen und fallbasiertes Schließen. Schweizerische Zeitschrift für Psychologie, 49:211–221, 1990.

Yokoyama:1990:OOC

[75] Tankanori Yokoyama. An objectoriented and constraint-based knowledge representation system for design object modeling. In Proceedings of the Sixth Conference on Artificial Intelligence Applications, pages 146–150. IEEE, Los Alamitos, CA, 1990.

chandra:1991:EID

[76] D. Navin chandra. Exploration and Innovation in Design: towards a Computational Model. Symbolic Computation — Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-97481-4. 194 pp.

Cunis:1991:PB

[77] Roman Cunis, Andreas Günther, and H. Strecker, editors. Das PLAKON-Buch: ein Expertensystemkern für Planungs- und Konfigurierungsaufgaben in technischen Domänen. Number 266 in Informatik Fachberichte. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-53683-3. viii + 279 pp. LCCN T57.6.P5 1991.

Domeshek:1991:TCB

[78] Eric A. Domeshek and Janet L. Kolodner. Toward a case-based aid for conceptual design. International Journal of Expert Systems, 4:201–220, 1991.

Fensel:1991:KKA

[79] Dieter Fensel, Jürgen Angele, and Dieter Landes. KARL: a knowledge acquisition and representation language. In J. C. Rault, editor, Proceedings of the 11th International Conference Expert Systems and their Applications, volume 1 (Tools, Techniques & Methods), pages 513–528. EC2, Nanterre, Avignon, 1991.

Frei:1991:PIR

[80] H. P. Frei, S. Meienberg, and P. Schäuble. The perils of interpreting recall and precision values. In Norbert Fuhr, editor, Information Retrieval: GI/GMD-Workshop Darmstadt, pages 1–10. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54618-9.

Gentner:1991:MFM

[81] Dedre Gentner and Kenneth D. Forbus. MAC/FAC: a model of similarity-based retrieval. In Proceedings of the 13th Cognitive Science Conference, pages 504– 509. Erlbaum, Hillsdale, Chicago, August 1991.

Harmelen:1991:FLK

[82] Frank van Harmelen and John Balder. (ML)^{2}: a formal language for KADS models of expertise. ES-PRIT Project P5248 KADS-II KADS-II/T1.2/PP/UvA/17/1.0, University of Amsterdam, Amsterdam, The Netherlands, November 1991.

Karbach:1991:WET

[83] Werner Karbach and Marc Linster. Wissensakquisition für Expertensysteme. Techniken, Modelle und Softwarewerkzeuge. Hanser, München, 1991. ISBN 3-446-15979-7. 196 pp.

Mandl:1991:KE

[84] Heinz Mandl, H. Gruber, and A. Renkl. Kontextualisierung von Expertise. Forschungsbericht 2, Ludwig-Maximilians-Universität. Lehrstuhl für Empirische Pädagogik und Pädagogische Psychologie, München, 1991.

Meinecke:1991:AKB

[85] Christoph Meinecke and Raimar G. Scherer. Architecture of a knowledgebased system for the detailing of reenforced concrete columns. In Gerhard N. Schmitt, editor, CAAD Futures ’91: Computer Aided Architectural Design Futures, pages 451–464. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1991. ISBN 3-528-08821-4.

Oess:1991:TQM

[86] Attila Oess. Total Quality Management: die ganzheitliche Qualitätsstrategie. Gabler Verlag, Wiesbaden, 2., erweiterte aufl. edition, 1991. ISBN 3-409-23622-8. 344 pp. LCCN ????

Pepperrell:1991:IMS

[87] Catherine A. Pepperrell. An Investigation of Methods for Similarity Searching in Databases of Three-Dimensional Chemical Structures. PhD thesis, University of Sheffield, Sheffield, UK, 1991. 320 pp. University Microfilms Order Number ADGDX-96527.

Richter:1991:SUC

[88] Michael M. Richter and Stefan Weß. Similarity, uncertainty and case-based reasoning in Patdex. In Robert S. Boyer, editor, Automated Reasoning, pages 249–265. Kluwer Academic Publishing, Dordrecht, 1991. Also appeared as SEKI-Report SR-91-01 (SFB), University of Kaiserslautern.

Wess:1991:PSA

[89] Stefan Weß. Patdex/2, ein System zum adaptiven, fallfokussierenden Lernen in technischen Diagnosesituationen. SEKI Working Paper SWP-91-01, Fachbereich Informatik, Universität Kaiserslautern, Kaiserslautern, Germany, December 1991. 162 pp.

Althoff:1992:FSE

[90] Klaus-Dieter Althoff, Stefan Weß, Brigitte Bartsch-Spörl, Frank Maurer, Dietmar Janetzko, and Angi Voß. Fallbasiertes Schließen in Expertensystemen. KI — Künstliche Intelligenz, 6(4): 14–21, 1992.

Balder:1992:FMK

[91] John Balder and Hans Akkermans. Formal Methods for Knowledge Modelling in the CommonKADS Methodology: a Compilation. Netherlands Enery Research Foundation, Petten, 1992. ISBN ???? 300 pp. LCCN ????

Bartsch-Sporl:1992:SIM

[92] Brigitte Bartsch-Spörl. A simple interpretation model for case-based reasoning. In Ch. Bauer and W. Karbach, editors, Interpretation models for KADS: Proceedings of 2nd KADS User Meeting, München, number 212 in GMD-Studien. GMD, Bonn, Germany, 1992.

Branskat:1992:FW

[93] Sonja Branskat. Fallbasierte Wissensakquisition. Arbeitspapiere 606, GMD, Bonn, Germany, 1992. 127 pp.

Bromme:1992:LEP

[94] Rainer Bromme. Der Lehrer als Experte: zur Psychologie des professionellen Wissens. Huber, Bern, 1992. ISBN 3-456-81991-9. 185 pp. LCCN ????

Christaller:1992:AWB

[95] Thomas Christaller, Franco di Primio, Uwe Schnepf, and Angi Voß. The AI-Workbench BABYLON — an Open and Portable Development Environment for Expert Systems. Academic Press, New York, USA, 1992. ISBN 0-12-174235-0. 474 pp.

Domeshek:1992:CBD

[96] Eric A. Domeshek and Janet L. Kolodner. A case-based design aid for architecture. In John S. Gero, editor, Artificial Intelligence in Design ’92, pages 497– 516. AID, Kluwer Academic Publishers, Dordrecht, Pittsburgh, 1992.

Durschke:1992:SBA

[97] Helge Dürschke and Wolfgang Oertel. Strukturierte Behandlung von Ähnlichkeiten bei der Nutzung großer Fallbibliotheken. In K. D. Althoff, S. Wess, B. Bartsch-Spörl, and D. Janetzko, editors, Proceedings: Ähnlichkeit von Fällen beim fallbasierten Schließen, Workshop, pages 42–46. Universität Kaiserslautern, FB Informatik, SEKI Working Paper SWP-92-11, 1992.

Esposito:1992:SDP

[98] Floriana Esposito, Giovanni Semeraro, and Donato Malerba. A syntactic distance for partially matching learned concepts against noisy structural object descriptions. International Journal of Expert Systems, 4:409–451, 1992.

Gauchel:1992:BMB

[99] J. Gauchel, S. van Wyk, R. R. Bhat, and Ludger Hovestadt. Building modeling based on concepts of autonomy. In J. Gero, editor, Artificial Intelligence in Design. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1992.

Gero:1992:AID

[100] John S. Gero, editor. Artificial Intelligence in Design: Papers from the 2nd Internat. Conf. on Artificial Intelligence in Design. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1799-8. 926 pp.

Grather:1992:AFS

[101] Wolfgang Gräther, Bernd Linowski, Michael Nowak, and Angi Voß. Ähnlichkeit von Fällen in Systemen des fallbasierten Schließens. In K. D. Althoff, S. Wess, B. Bartsch-Spörl, and D. Janetzko, editors, Proceedings: Ähnlichkeit von Fällen beim fallbasierten Schließen, Workshop, pages 51–56. Universität Kaiserslautern, FB Informatik, SEKI Working Paper SWP-92-11, 1992.

Greef:1992:ASU

[102] H. P. de Greef and Joost A. Breuker. Analysing system-user cooperation in KADS. Knowledge Acquisition, 4:89– 108, 1992.

Harmelen:1992:FLK

[103] Frank van Harmelen and John Balder. (ML)^{2}: a formal language for KADS models of expertise. Knowledge Acquisition, 4:127–161, 1992.

Harmelen:1992:KLR

[104] Frank van Harmelen, Bob Wielinga, Bert Bredeweg, Guus Schreiber, Werner Karbach, Martin Reinders, Angi Voß, Hans Akkermans, Brigitte Bartsch-Spörl, and Erik Vinkhuyzen. Knowledge level reflection. In Le Pate and L. Steels, editors, Enhancing the Knowledge Engineering Process — Contributions from ESPRIT, pages 175–204. Elsevier Publishing Company, Amsterdam, 1992.

Hinrichs:1992:PSO

[105] Thomas R. Hinrichs. Problem Solving in Open Worlds: a Case Study in Design. Lawrence Erlbaum Associates, Hillsdale, NJ, 1992. ISBN 0-8058-1228-8. ix + 228 pp. LCCN BF449 .H56 1992.

Hovestadt:1992:ADB

[106] Ludger Hovestadt. A4 — digitales bauen: Ein Modell für die weitgehende Computerunterstützung von Entwurf, Konstruktion und Betrieb von Gebäuden. In CAD im Architekturbüro: ACS-Kompendium ’92, pages 97–111. Architektenkammer Hessen, Wiesbaden, Wiesbaden, 1992.

Indurkhya:1992:MCI

[107] Bipin Indurkhya. Metaphor and Cognition: an Interactionist Approach. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1992. ISBN 0-7923-1687-8. 456 pp.

Jantke:1992:FCB

[108] Klaus P. Jantke. Formalizations in casebased reasoning. In K. D. Althoff, S. Wess, B. Bartsch-Spörl, and D. Janetzko, editors, Proceedings: Ähnlichkeit von Fällen beim fallbasierten Schließen, Workshop, pages 9–14. Universität Kaiserslautern, FB Informatik, SEKI Working Paper SWP-92-11, 1992.

Karbach:1992:ISA

[109] Werner Karbach and Angi Voß. An inference structure for assignment problems. In Bauer and W. Karbach, editors, Interpretation Models for KADS — Proc. 2nd KADS User Meeting, volume 212 of GMD-Studien, pages 323– 332. GMD, Bonn, Germany, 1992.

Kawamura:1992:LRF

[110] A. Kawamura, T. Minamikawa, K. Yura, A. Tanaka, T. Hayama, S. Masuda, and Y. Hidai. On-line recognition of freely handwritten Japanese characters using directional feature densities. In 11th IAPR International Conference on Pattern Recognition, volume II, pages 183– 186. Los Alamos, CA, IEEE Computer Society Press, The Hague, September 1992.

Knauff:1992:KAU

[111] Markus Knauff, R. Hörnig, Reinhold Rauh, and Gerhard Strube. Kontextsensitivität von Ähnlichkeitsurteilen: Empirie und Modellierung. In K. D. Althoff, S. Wess, B. Bartsch-Spörl, and D. Janetzko, editors, Proceedings: Ähnlichkeit von Fällen beim fallbasierten Schließen, Workshop, pages 75–78. Universität Kaiserslautern, FB Informatik, SEKI Working Paper SWP-92-11, 1992.

Komatsu:1992:RVC

[112] L. K. Komatsu. Recent views of conceptual structure. Psychological Bulletin, 112:500–528, 1992.

Lee:1992:SRS

[113] S.-Y. Lee and F.-J. Hsu. Spatial reasoning and similarity retrieval of images using 2D C-string knowledge representation. Pattern Recognition, 25:305–318, 1992. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Linster:1992:ARO

[114] Marc Linster, Werner Karbach, Angi Voß, and Jürgen Walther. An analysis of the role of operational modeling languages in the development of knowledgebased systems. In Mizoguchi, Motoda, Boose, Gaines, and Quinlan, editors, Proc. JKAW’92, pages 65–88. Hitachi Advanced Research Laboratories, ????, ????, 1992.

Linster:1992:KAB

[115] Marc Linster. Knowledge Acquisition Based on Explicit Methods of Problem-Solving. Diss., Universität Kaiserslautern, Kaiserslautern, Germany, 1992. 219 pp.

Minkoff:1992:FME

[116] M. Minkoff. The FERN model: an explanation with examples. Technical Report, Human Interface Technology Laboratory of the Washington Technology Center, University of Washington, Seattle, WA, USA, 1992.

Nemeth:1992:IAP

[117] Tibor Németh, Andreas Oberweis, Frank Schönthaler, and Wolffried Stucky. INCOME: Arbeitsplatz für den Programmentwurf interaktiver betrieblicher Informationssysteme. Bericht 251, Universität Karlsruhe (TH), Institut für angewandte Informatik und formale Beschreibungsverfahren, Karlsruhe, Germany, 1992.

Neubert:1992:CMC

[118] Susanne Neubert and Frank Maurer.

The conceptual model construction kit. In Proceedings of the Second European Conference on HyperText and Hypermedia. ACM Press, New York, NY 10036, USA, 1992.

Oberweis:1992:SCO

[119] Andreas Oberweis and Peter Sander.

The specification of complex object behaviour by high level Petri nets. Bericht 254, Universität Karlsruhe (TH), Institut für angewandte Informatik und formale Beschreibungsverfahren, Karlsruhe, Germany, 1992.

Rome:1992:WBG

[120] Erich Rome. Wahrnehmungspsychologie, Bilderkennung und der Grafikdesigner. TASSO-Report 36, GMD, Bonn, Germany, February 1992. 34 pp.

Schult:1992:FEM

[121] Th. J. Schult. Fallbasierte Expertensystemshells, Methoden und Werkzeuge. Forschungsbericht 89, Psychologisches Institut, Albert-Ludwigs-Universität Freiburg, December 1992.

Shepherd:1992:MSR

[122] Thomas S. Shepherd, William Uttal, Sriram Dayanand, and Robb Lovell. A method for shift, rotation, and scale invariant pattern recognition using the form and arrangement of patternspecific features. Pattern Recognition, 25:343–356, 1992. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Sycara:1992:CCB

[123] K. Sycara, D. Navin chandra, R. Guttal, J. Koning, and S. Narasimhan. CADET: a case-based synthesis tool for engineering design. International Journal of Expert Systems, 4:157–188, 1992.

Tuceryan:1992:SCE

[124] M. Tuceryan, A. K. Jain, and N. Ahuja. Supervised classification of early perceptual structure in dot patterns. In 11th IAPR International Conference on Pattern Recognition, volume II, pages 88– 91. IEEE Computer Society Press, Los Alamitos, CA, The Hague, September 1992.

Voss:1992:AKM

[125] Angi Voß and Werner Karbach. Ausführbare konzeptuelle Modelle. In R. Studer, editor, Informationssysteme und künstliche Intelligenz: Modellierung; 2. Workshop, Informatik-Fachberichte, pages 158–168. Springer, Berlin, Ulm, 1992.

Voss:1992:BCR

[126] Angi Voß and Werner Karbach. Building competent reflective systems. In A. Yonezawa and B. C. Smith, editors, Proc. IMSA ’92 of the International Workshop on Reflection and Metalevel Architecture, pages 154–158. Research Institute of Software Engineering (RISE) and Information-Technology Promotion Agency, Japan (IPA), ????, ????, 1992.

Voss:1992:MKM

[127] Angi Voß and Werner Karbach.

MODEL-K: making KADS run. In G. Schreiber, B. Wielinga, and J. Breuker, editors, KADS: a Principled Approach to Knowledge-Based System Development, pages 201–218. Academic Press, New York, USA, 1992.

Walther:1992:M

[128] Jürgen Walther, Angi Voß, Marc Linster, Thomas Hemmann, Hans Voß, and Werner Karbach. MoMo. Arbeitspapiere der GMD 658, GMD, Bonn, Germany, June 1992.

Wess:1992:FSU

[129] Stefan Weß, Jürgen Paulokat, and Klaus-Dieter Althoff, editors. Fallbasiertes Schließen — Ein Überblick – , Band I: Grundlagen. Universität Kaiserslautern, Fachbereich Informatik, Kaiserslautern, Germany, 1992. ISBN ???? 346 pp. LCCN ????

Wielinga:1992:KMA

[130] Bob J. Wielinga, A. Th. Schreiber, and Joost A. Breuker. KADS: a modelling approach to knowledge engineering. Knowledge Acquisition, 4:5–53, 1992.

Bardasz:1993:DCB

[131] Theodore Bardasz and Ibrahim Zeid.

DEJAVU: case-based reasoning for mechanical design. Artificial Intelligence for Engineering Design, Analysis and Manufacturing (AI EDAM), 72:111– 124, 1993.

Bhat:1993:CCB

[132] R. R. Bhat, J. Gauchel, and S. van Wyk. Communication in cooperative building design. In CAAD Futures 93: Proceedings of the Fifth International Conference on Computer-Aided Architectural Design Futures. North-Holland, Amsterdam, Pittsburgh, 1993.

Borner:1993:SSG

[133] Katy Börner. Structural similarity as guidance in case-based design. In S. Wess, K.-D. Althoff, and M. M. Richter, editors, Topics in Case-Based Reasoning: First European Workshop, EWCBR’93, pages 197–208. Springer, Berlin, Kaiserslautern, November 1993.

Buehrer:1993:ARC

[134] Daniel J. Buehrer and C. C. Chang. Application of a reciprocal confluence tree unit to similar-picture retrieval. In D. Abel and B. C. Ooi, editors, Advances in Spatial Databases: Proc., Third International Symposium SSD ’93, pages 437–442. Berlin: Springer, Singapur, June 1993.

Burkle:1993:PEO

[135] Ute Bürkle, Volker Weimer, and Heinz Züllighoven. Prototyping in einem objektorientierten Bankprojekt. In H. Züllighoven, W. Altmann, and E.-E. Doberkat, editors, Requirements Engineering ’93, Berichte des German Chapter of the ACM, pages 11–32. Teubner, Stuttgart, 1993.

Cawkell:1993:PQP

[136] A. E. Cawkell. Picture-queries and picture databases. Journal of Information Science, Principles and Practice, 19: 409–423, 1993. CODEN JISCDI. ISSN 0165-5515, 1352-7460.

Coulon:1993:IRR

[137] Carl-Helmut Coulon. Image retrieval without recognition. In M. M. Richter, S. Wess, K.-D. Althoff, and F. Maurer, editors, First European Workshop on Case-Based Reasoning (EWCBR’93), Posters and Presentations, volume 2 of SEKI-Report 93-12-2, pages 399– 402. Universität Kaiserslautern, Kaiserslautern, Germany, November 1993.

Domeshek:1993:FPL

[138] Eric A. Domeshek and Janet L. Kolodner. Finding the points of large cases. Artificial Intelligence for Engineering Design, Analysis and Manufacturing (AI EDAM), 7(2):87–96, 1993.

Fensel:1993:GSA

[139] Dieter Fensel, Jürgen Angele, Dieter Landes, and Rudi Studer. Giving structured analysis techniques a formal and operational semantics with KARL. In H. Züllighoven, W. Altmann, and E.-E. Doberkat, editors, Requirements Engineering ’93, Berichte des German Chapter of the ACM, pages 267–285. Teubner, Stuttgart, 1993.

Fidelak:1993:IPN

[140] Manfred Fidelak. Integritätsbedingungen in Petri-Netzen. PhD thesis, Universität Koblenz-Landau, Koblenz, December 1993. 181 pp.

Franke:1993:PMP

[141] W. Franke, U. Gutenbeil, W. Hasselbring, C. Pahl, H.-G. Sobottka, and B. Sucrow. Prototyping mit Mengen — der PROSET-Ansatz. In H. Züllighoven, W. Altmann, and E.- E. Doberkat, editors, Requirements Engineering ’93: Prototyping, Berichte des German Chapter of the ACM, pages 165–174. Stuttgart: Teubner, 1993.

Gramberg:1993:PFP

[142] Bernhard Gramberg. Der Prototyp als fertiges Produkt (oder Der Weg ist das Ziel). In H. Züllighoven, W. Altmann, and E.-E. Doberkat, editors, Requirements Engineering ’93, Berichte des German Chapter of the ACM, pages 33– 39. Teubner, Stuttgart, 1993.

Hoernig:1993:EIM

[143] R. Hoernig, R. Rauh, and Gerhard Strube. EVENTS-II: modeling event recognition. In G. Strube and K. F. Wender, editors, The Cognitive Psychology of Knowledge, pages 113–138. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1993.

Hovestadt:1993:ADBa

[144] Ludger Hovestadt. A4 — digitales bauen: Ein Modell für die weitgehende Computerunterstützung von Entwurf, Konstruktion und Betrieb von Gebäuden. PhD thesis, Institut für industrielle Bauproduktion der Universität Karlsruhe, Karlsruhe, Germany, 1993.

Hovestadt:1993:ADBb

[145] Ludger Hovestadt. A4 — digital building — extensive computer support for the design, construction, and management of buildings. In CAAD Futures ’93, Proceedings of the Fifth International Conference on Computer-Aided Architectural Design Futures, pages 405–422.

North-Holland, Amsterdam, Pittsburgh, June 1993.

Hua:1993:ECB

[146] K. Hua, Ian Smith, and Boi Faltings.

Exploring case-based design: CADRE. Artificial Intelligence for Engineering Design, Analysis and Manufacturing (AI EDAM), 7(2):135–144, 1993.

Karbach:1993:MKP

[147] Werner Karbach and Angi Voß.

MODEL-K for prototyping and strategic reasoning at the knowledge level. In Jean-Mark David, Jean-Paul Krivine, and Reid Simmons, editors, Second Generation Expert Systems, pages 721–745. Springer, Berlin, Berlin, 1993.

Knauff:1993:SAC

[148] Markus Knauff and Christoph Schlieder.

Similarity assessment and case representation in case-based design. In M. M. Richter, S. Wess, K.-D. Althoff, and F. Maurer, editors, First European Workshop on Case-Based Reasoning (EWCBR’93), Posters and Presentations, volume 1 of SEKI-Report 93-12-1, pages 37–42. Universität Kaiserslautern, November 1993.

Luo:1993:HSD

[149] Qiangyi Luo, P. G. Hendry, and J. T.

Buchanan. Heuristic search for distributed constraint satisfaction problem. Technical Report KEG-6-93, Dept. of Computer Science, Univ. of Strathclyde, Strathclyde, Scotland, February 1993. 28 pp.

Maiden:1993:CBR

[150] N. A. M. Maiden. Case-based reasoning in complex design tasks. In M. M. Richter, S. Wess, K.-D. Althoff, and F. Maurer, editors, EWCBR preprints, pages 252–257. Universität Kaiserslautern, Kaiserslautern, 1993.

Manago:1993:IRC

[151] Michel Manago, Klaus-Dieter Althoff, Eric Auriol, R. Traphöner, Stefan Wess, N. Conruyt, and Frank Maurer. Induction and reasoning from cases. In Michael M. Richter et al., editors, EWCBR-93 (preprints), number SR-93-12 in SEKI-Report, pages 313– 318. Fachbereich Informatik, Universität Kaiserslautern, 1993.

Montazeri:1993:ACB

[152] M. Montazeri and A. E. Adam. Applications of case-based reasoning to the law, the problems of multiple case reasoning, and indexing. In M. M. Richter, S. Wess, K.-D. Althoff, and F. Maurer, editors, EWCBR preprints, pages 43–47. Universität Kaiserslautern, Kaiserslautern, 1993.

Morgenstern:1993:AIB

[153] Karina Morgenstern. Anpassung im Bauentwurf mittels aktiver autonomer Objekte. Master’s thesis, Universität Bonn, Bonn, Germany, 1993. 184 pp.

Niblack:1993:QPQ

[154] Wayne Niblack, R. Barber, W. Equitz, Myron Flickner, E. Glasman, Dragutin Petkovic, Peter Yanker, Christos Faloutsos, and G. Taubin. The QBIC project: querying images by content using color, texture and shape. Research Report, IBM Research Division, Almaden Research Center, San Jose, California, February 1993.

Nowak:1993:KRR

[155] Michael Nowak, Angi Voß, and Wolfgang Gräther. Komplexe räumliche Relationen im Bauentwurfsprozeß. In D. Hernandez, editor, Forschungsberichte KI, Proc. Workshop Hybride und integrierte Ansätze zur Raumrepräsentation und ihre Anwendung, Report FKI-185-93, pages 44–49. ????, ????, 1993.

Ousterhout:1993:ITT

[156] John K. Ousterhout. An Introduction to Tcl and Tk. Addison-Wesley, Reading, MA, USA, 1993. ISBN ???? LCCN ????

Perner:1993:CBR

[157] P. Perner. Case-based reasoning for image interpretation in non-destructive testing. In M. M. Richter, S. Wess, K.-D. Althoff, and F. Maurer, editors, EWCBR preprints, pages 403–409. Univ. of Kaiserslautern, Kaiserslautern, 1993.

Pu:1993:IIC

[158] Pearl Pu. Introduction: Issues in casebased design systems. Artificial Intelligence for Engineering Design, Analysis and Manufacturing (AI EDAM), 7:79– 85, 1993.

Reitbock:1993:MMI

[159] H. J. Reitböck. Mechanismen der Mustererkennung im Sehsystem. In O. Herzog, Th. Christaller, and D. Schütt, editors, Grundlagen und Anwendungen der Künstlichen Intelligenz, pages 90–106.

17. Fachtagung für Künstliche Intelligenz, Berlin, Springer, 1993.

Rome:1993:MMG

[160] Erich Rome. MAX, ein maschinelles Gestalt-Erkennungssystem. KI — Künstliche Intelligenz, 7(Sonderheft): 70–71, 1993.

Schreiber:1993:KPA

[161] Guus Schreiber, Bob J. Wielinga, and Joost A. Breuker. KADS: a Principled Approach to Knowledge-Based System Development. Academic Press, New York, USA, 1993. ISBN 0-12-629040-7. 457 pp.

Silverman:1993:EEC

[162] B. G. Silverman and G. Wenig. Engineering expert critics for cooperative systems. The Knowledge Engineering Review, 8:309–328, 1993. CODEN KEREE3. ISSN 0269-8889.

Steinert:1993:EZA

[163] Dagmar Steinert. Effiziente Zugriffsstrukturen für die ähnlichkeitsbasierte Vorauswahl von Fällen. Diplomarbeit, Lehrstuhl für Informatik VI (KI) der Universität Würzburg, Würzburg, Germany, 1993.

Voss:1993:MDP

[164] Angi Voß, Hans Voß, Jürgen Walther, and Thomas Hemmann. Model-driven prototyping — prototype-driven modeling for knowledge-based systems. In Züllighoven, Altmann, and Doberkat, editors, Proc. RE’93, pages 303–316. Teubner, Stuttgart, Bonn, 1993.

Wess:1993:IRS

[165] Stefan Wess, Klaus-Dieter Althoff, and G. Derwand. Improving the retrieval step in case-based reasoning. In Michael M. Richter et al., editors, EWCBR-93 (preprints), number SR-93-12 in SEKI-Report, pages 83–88. Fachbereich Informatik, Universität Kaiserslautern, 1993.

Yang:1993:KKI

[166] Soon-Ae Yang, Dave Robertson, and John Lee. KICS: a knowledge-intensive case-based reasoning system for statutory building regulations and case histories. In The Fourth International Conference on Artificial Intelligence and Law: Proceedings of the Conference, June 1993, Amsterdam, pages 254–263. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-606-9.

Aamodt:1994:CBR

[167] Agnar Aamodt and Enric Plaza. Casebased reasoning: foundational issues, methodological variations, and system approaches. AI Communications, 7(1): 39–59, 1994.

Bakhtari:1994:ATA

[168] Shirin Bakhtari. An aptitude test for the application of case-based methods in open world domains. In Pearl Pu and Mary Lou Maher, editors, AID-94 Workshop Case-Based Design Systems. ????, ????, 1994.

Bakhtari:1994:BGB

[169] Shirin Bakhtari and Brigitte Bartsch-Spörl. Bridging the gap between AI technology and design requirements. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design’94, pages 753–768. Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Barletta:1994:HIR

[170] Ralph Barletta. A hybrid indexing and retrieval strategy for advisory CBR systems built with ReMind. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 49–58. AcknoSoft Press, Paris, 1994.

Bartsch-Sporl:1994:GBA

[171] Brigitte Bartsch-Spörl and Elisabeth-Ch. Tammer. Graph-based approach to structural similarity. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 45–58. GMD, Bonn, Germany, 1994.

Bartsch-Sporl:1994:MIP

[172] Brigitte Bartsch-Spörl, Shirin Bakhtari, Angi Voß, and Jörg W. Schaaf. The most important prerequisite for sharing knowledge and data in design is having achieved a common understanding. In Mert Alberts, editor, AID-94 Workshop A Semantic Basis for Sharing of Knowledge and Data in Design. ETH Zürich, Lausanne, 1994.

Bartsch-Sporl:1994:RUC

[173] Brigitte Bartsch-Spörl. Representing and using cases in visualisation-oriented design domains. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 253–261. AcknoSoft Press, Paris, Chantilly, 1994.

Bartsch-Sporl:1994:WND

[174] Brigitte Bartsch-Spörl. We need design cases which support the visual imagination of the designer. In Pearl Pu and Mary Lou Maher, editors, AID-94 Workshop Case-Based Design Systems. ????, ????, 1994.

Bath:1994:SSF

[175] Peter A. Bath, Andrew R. Poirrette, Peter Willett, and Frank H. Allen. Similarity searching in files of threedimensional chemical structures: comparison of fragment-based measures of shape similarity. Journal of Chemical Information and Computer Sciences, 34: 141–147, 1994. CODEN JCISD8. ISSN 0095-2338.

Bhatta:1994:IAD

[176] S. Bhatta, Ashok K. Goel, and S. Prabhakar. Innovation in analogical design: a model-based approach. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 57– 74. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2994-5.

Borner:1994:CBLa

[177] Katy Börner and Dietmar Janetzko.

Case-based learning for design. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 273–281. AcknoSoft Press, Paris, Chantilly, 1994.

Borner:1994:CBLb

[178] Katy Börner and Dietmar Janetzko.

Case-based learning for knowledgebased design support. In C. Nedellec and Y. Kodratoff, editors, Proceedings of the ECAI-94 Workshop on Integration of ML and KA, pages 13–17. Wiley, Chichester, Amsterdam, The Netherlands, 1994.

Borner:1994:TBA

[179] Katy Börner. Term-based approach to structural similarity as guidance for adaptation. In Angi Voß, editor, Similarity concepts and retrieval methods, volume 13 of Fabel-Reports, pages 59–72. GMD, Bonn, Germany, 1994.

Borner:1994:TFC

[180] Katy Börner. Towards formalizations in case-based reasoning for synthesis. In David W. Aha, editor, AAAI-94 Workshop on Case-Based Reasoning, Technical Report WS-94-01, pages 177–181. AAAI Press, Menlo Park, Seattle, Washington, 1994.

Bredeweg:1994:MBD

[181] Bert Bredeweg. Model-based diagnosis and prediction of behaviour. In Joost Breuker and Walter Van de Velde, editors, CommonKADS Library for Expertise Modelling, chapter 6, pages 121–153. IOS Press, Amsterdam, The Netherlands, 1994. ISBN 90-5199-164-9.

Breuker:1994:CLE

[182] Joost A. Breuker and Walter Van de Velde. CommonKADS Library for Expertise Modelling: Reusable Problem Solving Components. IOS Press, Amsterdam, The Netherlands, 1994. ISBN 90-5199-164-9. 359 pp.

Bures:1994:STD

[183] Mark G. Bures, Yvonne C. Martin, and Peter Willett. Searching techniques for databases of three-dimensional chemical structures. In Ernest L. Eliel and Samuel H. Wilen, editors, Topics in Stereochemistry, vol. 21, pages 467–511. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1994.

Clayton:1994:IOM

[184] M. J. Clayton, R. Fruchter, H. Krawinkler, and P. Teicholz. Interpretation objects for multi-disciplinary design. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 573–590. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994.

Coulon:1994:CFT

[185] Carl-Helmut Coulon and Ralf Steffens.

Comparing fragments by their images. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 36–44. GMD, Bonn, Germany, 1994.

Cross:1994:FIR

[186] Valerie Cross. Fuzzy information retrieval. Journal of Intelligent Information Systems, ??:29–56, 1994.

Dave:1994:CBD

[187] B. Dave, G. Schmitt, Boi Faltings, and Ian Smith. Case based design in architecture. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 145–162. AID, Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Domeshek:1994:DTK

[188] Eric A. Domeshek, Janet L. Kolodner, and C. M. Zimring. The design of a tool kit for case-based design aids. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 109–126. AID, Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Durschke:1994:HID

[189] Helge Dürschke. How to integrate different similarity concepts — a formal approach. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 105–108. GMD, Bonn, Germany, 1994.

Eltz:1994:SMI

[190] U. Eltz. System zum Monitoring von Inferenzprozessen. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1994.

Enzinger:1994:PS

[191] Andrea Enzinger, Frank Puppe, and Gerhard Strube. Problemlösen ohne Suchen? KI — Künstliche Intelligenz, 8(1):73–81, 1994.

Faloutsos:1994:EEQ

[192] Christos Faloutsos, R. Barber, Myron Flickner, Jim Hafner, Wayne Niblack, Dragutin Petkovic, and W. Equitz. Efficient and effective querying by image content. Journal of Intelligent Information Systems, 3:231–262, 1994.

Fensel:1994:CLW

[193] Dieter Fensel and Frank van Harmelen. A comparison of languages which operationalize and formalize KADS models of expertise. The Knowledge Engineering Review, 9:105–146, 1994. CODEN KEREE3. ISSN 0269-8889.

Frank:1994:CDG

[194] Steven Frank. Cataloging digital geographic data in the information infrastructure: a literature and technology review. Information Processing and Management, 30:587–606, 1994. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Giretti:1994:IAA

[195] A. Giretti, L. Spalazzi, and M. Lemma.

A.S.A. an interactive assistant to architectural design. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 93–108. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994.

Goldstone:1994:RSC

[196] Robert L. Goldstone. The role of similarity in categorization: providing a groundwork. Cognition, 52:125–157, 1994. CODEN CGTNAU. ISSN 0010-0277 (print), 1873-7838 (electronic).

Grather:1994:CDB

[197] Wolfgang Gräther. Computing distances between attribute-value representations in an associative memory. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 12–25. GMD, Bonn, Germany, 1994.

Gross:1994:UDA

[198] M. Gross, C. M. Zimring, and E. Do.

Using diagrams to access a case base of architectural designs. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 129–144. AID, Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Haigh:1994:GSM

[199] Karen Z. Haigh and Jonathan R.

Shewchuk. Geometric similarity metrics for case-based reasoning. In Case-Based Reasoning: Papers from the 1994 AAAI Workshop, Seattle, Aug. 1994, pages 157–163. AAAI Press, Menlo Park, CA, 1994. ISBN 0-929280-71-7.

Hemmann:1994:HSD

[200] Thomas Hemmann. The home stereo diagnosis revisited. In Hans Voß and Rudi Studer, editors, Proceedings 4th KADS Meeting, volume 832 of Arbeitspapiere der GMD. GMD, Sankt Augustin, Sankt Augustin, March 1994.

Hovestadt:1994:DFB

[201] Ludger Hovestadt and Barbara Schmidt-Belz. The domain of FABEL: building design. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 6–11. GMD, Bonn, Germany, 1994.

Hunt:1994:HCB

[202] John Hunt and Roger Miles. Hybrid case-based reasoning. The Knowledge Engineering Review, 9:383–397, 1994. CODEN KEREE3. ISSN 0269-8889.

Jaczynski:1994:FLR

[203] Michel Jaczynski and Brigitte Trousse. Fuzzy logic for the retrieval step of a case-based reasoner. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 313–321. AcknoSoft Press, Paris, 1994.

Janetzko:1994:MIS

[204] Dietmar Janetzko. Metaphorische Invention und semantische Innovation. Sprache & Kognition, 2:75–89, 1994.

Jantke:1994:NCS

[205] Klaus P. Jantke. Nonstandard concepts of similarity in case-based reasoning. In Hans-Hermann Bock, Wolfgang Lenski, and Michael M. Richter, editors, Information Systems and Data Analysis: Prospects — Foundations — Applications, Proceedings of the 17th Annual Conference of the GfKl, Univ. of Kaiserslautern, 1993, pages 28–43. Springer, Berlin, Kaiserslautern, 1994.

Kettler:1994:ECB

[206] Brian P. Kettler and James A. Hendler.

Evaluating a case-based planning system. In Case-Based Reasoning: Papers from the 1994 AAAI Workshop, Seattle, Aug. 1994, pages 157–163. AAAI Press, Menlo Park, CA, 1994. ISBN 0-929280-71-7.

Kindelberger:1994:EFS

[207] A. Kindelberger. Entwicklung eines fallbasierten Systems für die GO-Domäne. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1994.

Knauff:1994:DGC

[208] Markus Knauff and Christoph Schlieder.

Dynamic grouping: case reinterpretation as a foundation of knowledgeintensive similarity assessment. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 85–104. GMD, Bonn, Germany, 1994.

Knauff:1994:I

[209] Markus Knauff. Introduction. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 1–5. GMD, Bonn, Germany, 1994.

Knauff:1994:MGP

[210] Markus Knauff. Modeling gestalt principles as guidance for dynamic grouping. In Proceedings of the First Workshop on Visual Computing, pages 113– 121. Fraunhofer-IGD, Darmstadt, 1994.

Layne:1994:SII

[211] Sara Shatford Layne. Some issues in the indexing of images. Journal of the American Society for Information Science, 45:583–588, 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Linowski:1994:CDB

[212] Bernd Linowski. Computing distances between attribute-value representations in a flat memory. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 26–35. GMD, Bonn, Germany, 1994.

Luo:1994:MUI

[213] Qiangyi Luo, P. G. Hendry, and J. T. Buchanan. Methods used to implement an integrated distributed scheduler. In S. M. Deen, editor, Proceedings of the CKBS-SIG Workshop, Sept. 1993, pages 47–60. DAKE Centre, University of Keele, Keele, UK, 1994.

Marir:1994:CBR

[214] Farhi Marir and Ian Watson. Case-based reasoning: a categorized bibliography. The Knowledge Engineering Review, 9: 355–381, 1994. CODEN KEREE3. ISSN 0269-8889.

Oxman:1994:PDC

[215] Rivka E. Oxman. Precedents in design: a computational model for the organisation of precedent knowledge. Design Studies, 17(2):117–134, 1994.

Petrakis:1994:SSL

[216] Euripides G. M. Petrakis and Christos Faloutsos. Similarity searching in large image databases. Technical Report CS-TR-3388, University of Maryland, College Park, MD, USA, 1994. 33 pp.

Pu:1994:FCA

[217] Pearl Pu and Lisa Purvis. Formalizing case adaptation in a case-based design system. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design ’94, pages 77–91. AID, Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Puppe:1994:WBK

[218] Frank Puppe, Karsten Poeck, Ute Gappa, Stefan Bamberger, and Klaus Goos. Wiederverwendbare Bausteine für eine konfigurierbare Diagnostik-Shell. KI — Künstliche Intelligenz, 8(2):13–18, 1994.

Ricci:1994:CPE

[219] Francesco Ricci, S. Mam, P. Marti, V. Normand, and P. Olmo. CHARADE: a platform for emergencies management systems. Technical Report 9404-07, Istituto per la Ricerca Scientifica e Tecnologica, Povo, 1994. 14 pp.

Schaaf:1994:DGC

[220] Jörg W. Schaaf. Detecting gestalts in CAD-plans to be used as indices. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 73–84. GMD, Bonn, Germany, 1994.

Schaaf:1994:GCP

[221] Jörg W. Schaaf. Gestalts in CAD-plans, analysis of a similarity concept. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design’94, pages 437–446. Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Schult:1994:CBE

[222] Thomas Josef Schult and Dietmar Janetzko. Case-based expert system shells: first and second generation. In Jay Liebowitz, editor, Moving Towards Expert Systems Globally in the 21st Century, pages 306–310. Cognizant Communication Corporation, Elmsford, NY, Ney York, 1994.

Voss:1994:RSLa

[223] Angi Voß. Retrieval of similar layouts. In Angi Voß, editor, Similarity Concepts and Retrieval Methods, volume 13 of Fabel-Reports, pages 109–124. GMD, Bonn, Germany, 1994.

Voss:1994:RSLb

[224] Angi Voß, Carl-Helmut Coulon, Wolfgang Gräther, Bernd Linowski, Jörg W. Schaaf, Brigitte Bartsch-Spörl, Katy Börner, Elisabeth-Ch. Tammer, Helge Dürschke, and Markus Knauff. Retrieval of similar layouts — about a very hybrid approach in FABEL. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design’94, pages 625–640. Kluwer Academic Publishers, Dordrecht, Lausanne, 1994.

Voss:1994:RSLc

[225] Angi Voß. Retrieval of similar layouts in FABEL. In ikm: Internationales Kolloquium über Anwendungen der Informatik und der Mathematik in Architektur und Bauwesen, Abstracts, pages 163–168. Hochschule für Architektur und Bauwesen, Weimar, Weimar, March 1994.

Walther:1994:OAM

[226] Jürgen Walther, Wolfgang Gräther, Wolfgang Oertel, Barbara Schmidt-Belz, and Angi Voß. An open architecture for multiple case retrieval methods. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 373–380. AcknoSoft Press, Paris, Chantilly, November 1994.

Wolf:1994:EAB

[227] T. Wolf. Entwicklung allgemeiner Bausteine zur Erstellung fallbasierter Systeme. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1994.

Adami:1995:AAA

[228] Parivash Adami. Adaptation by active autonomous objects (AAAO). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 46–50. GMD, Bonn, Germany, June 1995.

Althoff:1995:RIC

[229] Klaus-Dieter Althoff, Eric Auriol, Ralph Barletta, and Michel Manago. A Review of Industrial Case-Based Reasoning Tools. AI Intelligence, Oxford, 1995. ISBN 1-898804-01-X. 148 pp.

Angele:1995:MIK

[230] Jürgen Angele, Dieter Fensel, Dieter Landes, and Susanne Neubert. Modellbasiertes und inkrementelles Knowledge Engineering: der MIKE-Ansatz. KI — Künstliche Intelligenz, 9(1):16–21, 1995.

Bakhtari:1995:DAA

[231] Shirin Bakhtari and Wolfgang Oertel. DOM: an active assistance system for architectural and engineering design. In Milton Tan and Robert Teh, editors, The Global Design Studio — Proceedings of the 6th International Conference on Computer-Aided Architectural Design Futures, pages 153–162. Centre for Advanced Studies in Architecture, National University of Singapore, Singapore, 1995.

Bakhtari:1995:QAD

[232] Shirin Bakhtari and Wolfgang Oertel. Quality assessment of design cases within the DOM environment. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Wess, editors, Proceedings of the 3rd German CBR Workshop, pages 106–113. LSA-95-02 University of Kaiserslautern, Kaiserslautern, 1995.

Bartsch-Sporl:1995:ASD

[233] Brigitte Bartsch-Spörl, Shirin Bakhtari, and Wolfgang Oertel. Assessment supported by a domain ontology (DOM). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 23–30. GMD, Bonn, Germany, June 1995.

Bartsch-Sporl:1995:KMI

[234] Brigitte Bartsch-Spörl. KI-Methoden für innovative Design-Domänen. In Michael Richter and Frank Maurer, editors, Expertensysteme 95, pages 137– 151. infix, Sankt Augustin, Kaiserslautern, 1995.

Bergmann:1995:BRA

[235] Ralph Bergmann and Wolfgang Wilke. Building and refining abstract planning cases by change of representation language. Journal of Artificial Intelligence Research, 3:53–118, 1995.

Bhat:1995:CAU

[236] Raghu R. Bhat. Case adaptation using agents (AgentEx). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 51–56. GMD, Bonn, Germany, June 1995.

Borner:1995:ALD

[237] Katy Börner and Roland Faßauer. Analogical layout design (Syn*). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 59–68. GMD, Bonn, Germany, June 1995.

Borner:1995:CA

[238] Katy Börner. Conceptual analogy. In D. W. Aha and A. Ram, editors, AAAI 1995 Fall Symposium Series: Adaptation of Knowledge for Reuse, Technical Report FS-95-04, pages 5–11. AAAI Press, Menlo Park, Boston, MA, 1995.

Borner:1995:CM

[239] Katy Börner and Gerhard Strube. Comparison of the modules. In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 83–89. GMD, Bonn, Germany, June 1995.

Borner:1995:DSI

[240] Katy Börner. Design support in industrial building design. In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 3–12. GMD, Bonn, Germany, June 1995.

Borner:1995:IAC

[241] Katy Börner. Interactive, adaptive, computer aided design. In Milton Tan and Robert Teh, editors, The Global Design Studio — Proceedings of the 6th International Conference on Computer-Aided Architectural Design Futures, pages 627–634. Centre for Advanced Studies in Architecture, National University of Singapore, Singapore, 1995.

Borner:1995:MTO

[242] Katy Börner and Dietmar Janetzko. A methodology to task-oriented knowledge acquisition and system design. In Barbara Schmidt-Belz, editor, Scenario of FABEL Prototype 3 Supporting Architectural Design, volume 40 of Fabel-Report. GMD, Bonn, Germany, 1995.

Borner:1995:SAC

[243] Katy Börner and Dietmar Janetzko. System architecture for computer-aided building engineering. In Peter Jan Pahl and Heinrich Werner, editors, 6th International Conference on Computing in Civil and Building Engineering, pages 385–390. Balkema, Rotterdam, Berlin, 1995.

Branting:1995:SCB

[244] L. Karl Branting and David W. Aha. Stratified case-based reasoning: reusing hierarchical problem solving episodes. In IJCAI-95: Proceedings of the 14th International Joint Conference on Artificial Intelligence, pages 384–390. Morgan Kaufmann, San Mateo, CA, Montreal, August 1995.

Christaller:1995:PIK

[245] Thomas Christaller. Programmierstil im Klassenformat. Schweizerische Technische Zeitschrift, 9:26–29, 1995.

Coulon:1995:AIRa

[246] Carl-Helmut Coulon. Automatic indexing, retrieval and reuse of topologies in architectural layouts. In Milton Tan and Robert Teh, editors, The Global Design Studio — Proceedings of the 6th International Conference on Computer-Aided Architectural Design Futures, pages 577– 586. Centre for Advanced Studies in Architecture, National University of Singapore, Singapore, 1995.

Coulon:1995:AIRb

[247] Carl-Helmut Coulon. Automatic indexing, retrieval and reuse of topologies in complex designs. In Peter Jan Pahl and Heinrich Werner, editors, Proceeding of the Sixth International Conference on Computing in Civil and Building Engineering, pages 749–754. Balkema, Rotterdam, Berlin, 1995.

Coulon:1995:BDL

[248] Carl-Helmut Coulon. Benutzung von Datensammlungen zum Layout. In Michael Richter and Frank Maurer, editors, Expertensysteme 95, pages 170– 186. infix, Sankt Augustin, Kaiserslautern, 1995. Proceedings of the German Conference on Expert Systems.

Coulon:1995:GGT

[249] Carl-Helmut Coulon. General geometric and topologic retrieval and adaptation (TOPO). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 33–45. GMD, Bonn, Germany, June 1995.

Deters:1995:MSC

[250] Ralph Deters. Multidimensional structuring of a case-memory. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Wess, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 21– 29. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, 1995.

Enser:1995:PIR

[251] P. G. B. Enser. Pictorial information retrieval. Journal of Documentation, 51:126–170, 1995. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Grather:1995:ECP

[252] Wolfgang Gräther. Exact and correct placement of pipes (ANOPLA). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 69–77. GMD, Bonn, Germany, June 1995.

Heene:1995:AGL

[253] S. Heene. Automatische Generierung von Lageplänen. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1995.

Hinkle:1995:ACB

[254] David Hinkle and C. Toomey. Applying case-based reasoning to manufacturing. The AI Magazine, 16(1):65–73, 1995. CODEN AIMAEK. ISSN 0738-4602.

Hovestadt:1995:PMB

[255] Ludger Hovestadt and Barbara Schmidt-Belz. PM5 — a model of building design. In Barbara Schmidt-Belz, editor, Scenario of FABEL Prototype 3 Supporting Architectural Design, volume 40 of Fabel-Report. GMD, Bonn, Germany,

Hunt:1995:TIA

[256] John Hunt and Roger Miles. Toward an intelligent architectural design aid. Expert Systems, 12:209–218, 1995.

Janetzko:1995:ARD

[257] Dietmar Janetzko and Oliver Jäschke.

Assessing realizations of design tasks (CheckL & CheckUp). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 15–22. GMD, Bonn, Germany, June 1995.

Janetzko:1995:LHT

[258] Dietmar Janetzko and Thomas Josef Schult. LanPlan: a hypermedia tutor using multiple representations. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Weß, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 81–86. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, 1995.

Janetzko:1995:UCA

[259] Dietmar Janetzko and Oliver Jäschke.

Using cases to acquire knowledge for quality control in design. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Weß, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 114–122. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, 1995.

Janetzko:1995:UCQ

[260] Dietmar Janetzko and Oliver Jäschke.

Using cases for quality control in design. In Peter Jan Pahl and Heinrich Werner, editors, Proceeding of the Sixth International Conference on Computing in Civil and Building Engineering, pages 187–192. Balkema, Rotterdam, Berlin, 1995.

Jaschke:1995:OUR

[261] Oliver Jäschke and Dietmar Janetzko. Operators used in routine design (Roude). In Katy Börner, editor, Modules for Design Support, volume 35 of Fabel-Report, pages 78–80. GMD, Bonn, Germany, June 1995.

Jensen:1995:CPN

[262] Kurt Jensen. Coloured Petri Nets: Basic Concepts, Analysis Methods and Practical Use, vol. 2. EATCS Monographs on Theoretical Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58276-2. 174 pp.

Lenz:1995:RS

[263] Mario Lenz and Hans-Dieter Burkhard. Retrieval ohne Suche. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Wess, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 1– 10. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, Germany, 1995.

Loftus:1995:DSE

[264] C. W. Loftus, E. M. Sherratt, R. J. Gautier, P. A. M. Grandi, D. E. Price, and M. D. Tedd. Distributed Software Engineering. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-342676-9. 260 pp.

Maser:1995:WAS

[265] M. Mäser. Wissensbasierte automatische Steuerung und Regelung hochautomatisierter Gebäude. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1995.

Netten:1995:EAC

[266] Bart D. Netten, R. A. Vingerhoeds, and H. Koppelaar. Expert assisted conceptual design: an application to fibre reinforced composite panels. In R. Oxman, M. Bax, and H. Achten, editors, Design Research in the Netherlands, pages 125– 139. Faculty of Architecture, Planning, and Building Design Science, Eindhoven University, Eindhoven, 1995.

Petzoldt:1995:PPD

[267] Th. Petzoldt, P. Kielb, Uwe Petersohn, and N. Schütze. Prognose der Phytoplankton-Dynamik: Anwendung fallbasierter Methoden auf einen dynamischen ökologischen Prozeß. In M. Hilty and A. Jäschke, editors, Informatik für den Umweltschutz, pages 259– 267. Metropolis-Verlag, Marburg, 1995.

Schaaf:1995:MSN

[268] Jörg W. Schaaf and Angi Voß. Ein Modell zur Suche nützlicher Fälle im fallbasierten Schließen. In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Weß, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 11–20. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, 1995.

Schaaf:1995:RSL

[269] Jörg W. Schaaf and Angi Voß. Retrieval of similar layouts in FABEL using AspecT. In Milton Tan and Robert Teh, editors, The Global Design Studio — Proceedings of the 6th International Conference on Computer-Aided Architectural Design Futures, pages 617–626. Centre for Advanced Studies in Architecture, National University of Singapore, 1995.

Schaaf:1995:WAW

[270] Jörg W. Schaaf. Wechselnde Aspekte oder: Wie strukturiert man eine Fallbasis, ohne zu wissen wonach? In Brigitte Bartsch-Spörl, Dietmar Janetzko, and Stefan Weß, editors, Fallbasiertes Schließen — Grundlagen und Anwendungen, LSA-Report 95-02, pages 11–20. Zentrum für Lernende Systeme und Anwendungen, Fachbereich Informatik, Universität Kaiserslautern, 1995.

Schmidt-Belz:1995:PFI

[271] Barbara Schmidt-Belz. Profiles of the FABEL IDEA 3 tools. In Barbara Schmidt-Belz, editor, Scenario of FA-BEL Prototype 3 Supporting Architectural Design, volume 40 of Fabel-Report. GMD, Bonn, Germany, 1995.

Schutze:1995:EIS

[272] N. Schütze. Entwurf und Implementierung eines Systems zum fallbasierten Schließen für die Prognose von Algenmassenentwicklungen in Standgewässern. Diplomarbeit, Fakultät für Informatik, TU Dresden, Dresden, Germany, 1995.

Veloso:1995:CBR

[273] Manuela M. Veloso and Agnar Aamodt, editors. Case Based Reasoning Research and Development: First International Conference, ICCBR-95. Number 1010 in Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60598-3. 576 pp.

Voss:1995:CA

[274] Angi Voß and Werner Karbach. Competence assessment. In Peter Hoschka, editor, Computers as Assistants — a New Generation of Support Systems, chapter 2.1, pages 17–28. Lawrence Erlbaum Associates, Hillsdale, Cambridge MA, 1995. ISBN 0-8058-2188-0.

Voss:1995:CBR

[275] Angi Voß. Case-based reasoning in building design: problems of case elicitation and retrieval. International Journal of Construction Information Technology, 28(4):49–62, 1995.

Voss:1995:IDA

[276] Angi Voß, Wolfgang Gräther, and Jörg W. Schaaf. Intelligent design assistant. In Peter Hoschka, editor, Computers as Assistants — A New Generation of Support Systems, chapter 3.5, pages 142–155. Lawrence Erlbaum Associates, Hillsdale, Cambridge MA, 1995. ISBN 0-8058-2188-0.

Voss:1995:RTD

[277] Angi Voß and Brigitte Bartsch-Spörl.

Retrieval of technical design documents for reuse. In Peter Pahl and Heinrich Werners, editors, Computing in Civil and Building Engineering, pages 923– 929. Balkema, Rotterdam, Berlin, 1995.

Voss:1995:TMA

[278] Angi Voß. Towards mindful artifacts. In John S. Gero and Fay Sudweeks, editors, Fourth Workshop on Research Directions for Artificial Intelligence in Design, pages 87–90. University of Sydney, University of Twente, Enschede The Netherlands, January 1995.

Yang:1995:UCB

[279] Soon-Ae Yang, Dave Robertson, and John Lee. Use of case-based reasoning in the domain of building regulations. In J.-P. Haton, M. Keane, and M. Manago, editors, Advances in Case-Based Reasoning: Second European Workshop, EWCBR-94, pages 292–306. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60364-6.

Althoff:1996:DSC

[280] Klaus-Dieter Althoff and Brigitte Bartsch-Spörl. Decision support for case-based applications. Wirtschaftsinformatik, 38:8–16, 1996. CODEN WI-INE9. ISSN 0937-6429.

Althoff:1996:RCB

[281] Klaus-Dieter Althoff and Agnar Aamodt. Relating case-based problem solving and learning methods to task and domain characteristics: towards an analytic framework. AI Communications, 9:109– 116, 1996.

Auriol:1996:CFC

[282] Eric Auriol, Michel Manago, and Jérôme Guiot-Dorel. CASSIOPÉE: Fehlerdiagnose von CFM 56-3 Triebwerken für Boeing 737 Flugzeuge. KI — Künstliche Intelligenz, 10(1):47–53, 1996.

Bakhtari:1996:DAA

[283] Shirin Bakhtari, Brigitte Bartsch-Spörl, and Wolfgang Oertel. DOM-ARCADE: assistance services for construction, evaluation, and adaptation of design layouts.

In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design’96, pages 681–699. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1996. ISBN 0-7923-4080-9.

Bartsch-Sporl:1996:HMC

[284] Brigitte Bartsch-Spörl. How to make CBR systems work in practice. In H.-D. Burkhard and M. Lenz, editors, Fourth German Workshop on Case-Based Reasoning: System Development and Evaluation, number 55 in Informatik-Berichte, pages 36–42. Humboldt-Universität, Berlin, 1996.

Bartsch-Sporl:1996:SSB

[285] Brigitte Bartsch-Spörl and Shirin Bakhtari. A support system for building design — experiences and convictions from the FABEL project. In John Sharpe, editor, AI System Support for Conceptual Design: Proceedings of the 1995 Lancaster International Workshop on Engineering Design, pages 279–297. Springer, Berlin, London, 1996.

Borner:1996:CAC

[286] Katy Börner, Heiko Wode, and Roland Faßauer. Conceptual analogy: conceptual clustering for informed and efficient analogical reasoning. In H.-D. Burkhard and M. Lenz, editors, Fourth German Workshop on Case-Based Reasoning: System Development and Evaluation, number 55 in Informatik-Berichte, pages 78–86. Humboldt-Universität, Berlin, 1996.

Burkhard:1996:CRN

[287] Hans-Dieter Burkhard and Mario Lenz.

Case retrieval nets: basic ideas and extensions. In H.-D. Burkhard and M. Lenz, editors, Fourth German Workshop on Case-Based Reasoning: System Development and Evaluation, number 55 in Informatik-Berichte, pages 103–110. Humboldt-Universität, Berlin, 1996.

Coulon:1996:RAI

[288] Carl-Helmut Coulon. Die Rolle des Anpassungswissens im CBR. In H. Czap, P. Jaenecke, and H. P. Ohly, editors, Analogie in der Wissensrepräsentation: Case-Based Reasoning und räumliche Modelle, 4. Tagung der deutschen Sektion der Internationalen Gesellschaft für Wissensrepräsentation, Okt. 1995, Trier, pages 34–41. Indeks Verlag, Frankfurt, 1996. ISBN 3-88672-405-0.

Goos:1996:FKM

[289] Klaus Goos. Fallbasiertes Klassifizieren: Methoden, Integration und Evaluation. Number 127 in Dissertationen zur Künstlichen Intelligenz. infix, Sankt Augustin, 1996. ISBN 3-89601-127-8. 180 pp. Also: Würzburg, Universität, Diss. 1995.

Hemmann:1996:ROK

[290] Thomas Hemmann. Reuse-Oriented Knowledge Engineering with MoMo: an Inquiry. PhD thesis, Universität Bielefeld, Bielefeld, June 1996. Also: R. Oldenbourg, München (Berichte der GMD, 267), ISBN 3-486-24002-1.

Leake:1996:CBR

[291] David B. Leake, editor. Case-Based Reasoning: Experiences, Lessons, and Future Directions. AAAI Press, Menlo Park, CA, 1996. ISBN 0-262-62110-X. 420 pp.

Lenz:1996:FDE

[292] Mario Lenz, Eric Auriol, Hans-Dieter Burkhard, Michel Manago, and Petra Pirk. Fallbasierte Diagnostik und Entscheidungsunterstützung. KI — Künstliche Intelligenz, 10(1):16–21, 1996.

Maher:1996:CBR

[293] Mary Lou Maher, M. Bala Balachandran, and Dong Mei Zhang. Case-Based Reasoning in Design. Lawrence Erlbaum Associates, Mahwah, NJ, 1996. ISBN 0-8058-1832-4. 246 pp.

Milosavljevic:1996:RA

[294] Aleksandar Milosavljević. Repeat analysis. In Nigel Spurr, editor, ICRF Handbook of Genome Analysis, chapter 13.4. Blackwell Science, Oxford, 1996. ISBN 0-632-03728-8.

Oertel:1996:IFR

[295] Wolfgang Oertel. Zur Integration von fall- und regelbasierten Verfahren im Entwicklungssystem FAENSY. In H.-D. Burkhard and M. Lenz, editors, Fourth German Workshop on Case-Based Reasoning: System Development and Evaluation, number 55 in Informatik-Berichte, pages 128–135. Humboldt-Universität, Berlin, March 1996.

Oertel:1996:IGK

[296] Wolfgang Oertel and Shirin Bakhtari.

Interaction of generic knowledge and cases in DOM. In Computing in Civil Engineering: Proceedings of the Third Congress. American Society of Civil Engineers, New York, Anaheim, USA, June

Oxman:1996:CBD

[297] Rivka Oxman and Angi Voß. Case-based design. AI Communications, 9:117–127, 1996. Also published in AID’96, Workshop Notes New Directions in Case-Based Design Systems, Stanford, June 1996, pp. 32-36.

Oxman:1996:FE

[298] Rivka Oxman and Angi Voß. Fallgestütztes Entwerfen. KI — Künstliche Intelligenz, 10(1):29–35, 1996.

Rissland:1996:BSL

[299] Edwina L. Rissland, David B. Skalak, and M. Timur Friedman. BankXX: supporting legal arguments through heuristic retrieval. Artificial Intelligence and Law, 4:1–71, 1996.

Schaaf:1996:FV

[300] Jörg W. Schaaf. Fischen und Versenken. In H. Czap, P. Jaenecke, and H. P. Ohly, editors, Analogie in der Wissensrepräsentation: Case-Based Reasoning und räumliche Modelle, 4. Tagung der deutschen Sektion der Internationalen Gesellschaft für Wissensrepräsentation, Okt. 1995, Trier, pages 42–49. Indeks Verlag, Frankfurt, 1996. ISBN 3-88672-405-0.

Schmidt-Belz:1996:HSC

[301] Barbara Schmidt-Belz, Angi Voß, Lars Emkes, and Carl-Helmut Coulon. How to support city planning using map interpretation techniques. In Harry Timmermans, editor, 3rd Design & Decision Support Systems in Architecture & Urban Planning Conference, Spa, Belgium, August 1996, volume 2, pages 228–245. Eindhoven University of Technology, Eindhoven, 1996. ISBN 90-681-065-5 (??invalid ISBN checksum??).

Schmidt-Belz:1996:SID

[302] Barbara Schmidt-Belz and Ludger Hovestadt. Scenario of an integrated design support for architects. Design Studies, 17:498–509, 1996.

Strube:1996:CCE

[303] Gerhard Strube, Dietmar Janetzko, and Markus Knauff. Cooperative construction of experts knowledge: The case of knowledge engineering. In P. B. Baltes and U. M. Staudinger, editors, Interactive Minds. Cambridge University Press, Cambridge, UK, 1996.

Tague-Sutcliffe:1996:SPE

[304] Jean M. Tague-Sutcliffe. Some perspectives on the evaluation of information retrieval systems. Journal of the American Society for Information Science, 47:1–3, 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voss:1996:F

[305] Angi Voß, Brigitte Bartsch-Spörl, Ludger Hovestadt, Klaus P. Jantke, Uwe Petersohn, and Gerhard Strube. FA-BEL. KI — Künstliche Intelligenz, 10 (3):70–76, 1996.

Voss:1996:HSC

[306] Angi Voß. How to solve complex problems with cases. Engineering Applications of Artificial Intelligence, 9:377– 384, 1996.

Voss:1996:SAT

[307] Angi Voß and Carl-Helmut Coulon.

Structural adaptation with TOPO. In Angi Voß, editor, ECAI 96, W[orkshop] 6, Adaptation in Case-Based Reasoning, pages 52–54. ECAI 96, 1996.

Voss:1996:SCA

[308] Angi Voß, Brigitte Bartsch-Spörl, and Rivka Oxman. A study of case adaptation systems. In John S. Gero and Fay Sudweeks, editors, Artificial Intelligence in Design’96, pages 173–189. Kluwer Academic Publishers, Dordrecht, Stanford, 1996. ISBN 0-7923-4080-9.

Wess:1996:ISC

[309] Stefan Wess. Intelligente Systeme für den Computer-Support: fallbasiertes Schließen in Help-Desk- und Call-Center-Anwendungen. Wirtschaftsinformatik, 38:23–31, 1996. CODEN WI-INE9. ISSN 0937-6429.

Woltering:1996:ESS

[310] Ansgar Woltering and Stefan Wess. Entscheidungsunterstützungssysteme und fallbasiertes Schließen: eine empirische Untersuchung im deutschsprachigen Raum. Wirtschaftsinformatik, 38:17–22, 1996. CODEN WIINE9. ISSN 0937-6429.

Maher:1997:IAC

[311] Mary Lou Maher and Pearl Pu. Issues and Applications of Case-Based Reasoning in Design. Lawrence Erlbaum Associates, Mahwah, NJ, 1997. ISBN 0-8058-2313-1. 342 pp.

Voss:1997:CDS

[312] Angi Voß. Case design specialists in FA-BEL. In Mary Lou Maher and Pearl Pu, editors, Issues and Applications of Case Based Reasoning to Design, pages 301–336. Lawrence Erlbaum Associates, Mahwah, NJ, 1997. ISBN 0-8058-2313-1.

Voss:1997:CRS

[313] Angi Voß. Case reusing systems — survey, framework and guidelines. The Knowledge Engineering Review, 12(1): 59–89, 1997. CODEN KEREE3. ISSN 0269-8889.

Budde:1983:APP

[314] Reinhard Budde, K. Kuhlenkamp, L. Mathiasen, and Heinz Züllighoven, editors. Approaches to Prototyping: Proc. of the Working Conf. on Prototyping. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 1984. ISBN 3-540-13490-5.

Hammond:1989:CBR

[315] Kristian J. Hammond, editor. Casebased reasoning: proceedings of a Workshop on Case-Based Reasoning: Holiday Inn, Pensacola Beach, Florida, May 31– June 2, 1989. DARPA, Morgan Kaufmann, San Mateo, CA, USA, May 1989. ISBN 1-55860-068-X. LCCN BF442 .W67 1989.

DARPA:1991:CBR

[316] Defense Advanced Research Projects Agency, editor. Case-based reasoning: proceedings of a Workshop held at the Madison Hotel, Washington, DC, May 8–10, 1991. Morgan Kaufmann, San Mateo, CA, USA, May 1991. ISBN 1-55860-199-6. LCCN BF442 .W67 1991.

Bareiss:1991:PCB

[317] Ray Bareiss, editor. Case-based reasoning: proceedings of a workshop held at the Madison Hotel, Washington, DC, May 8–10, 1991. Morgan Kaufmann, San Mateo, 1991. ISBN 1-55860-199-6. LCCN BF442 .W67 1991.

Zullighoven:1993:REP

[318] Heinz Züllighoven, W. Altmann, and E.- E. Doberkat, editors. Requirements Engineering ’93: Prototyping, number 41 in Berichte des German Chapter of the ACM. Teubner, Stuttgart, Germany; Leipzig, Germany, 1993. ISBN 3-519-02682-1.

Leake:1993:CBR

[319] David B. Leake, editor. Case-based reasoning: papers from the 1993 workshop, July 11–12, Washington, DC, Technical Report WS-93-01. AAAI, Menlo Park, CA, July 1993. ISBN ???? LCCN ????

Visser:1993:RDI

[320] Willemien Visser, editor. Reuse of designs: an interdisciplinary cognitive approach: proceedings of the Workshop of the Thirteenth International Joint Conference on Artificial Intelligence, Chambery, France, August 29, 1993. INRIA, Le Chesnay, France, August 1993. ISBN 2-7261-0815-6. LCCN ????

Wess:1994:TCR

[321] Stefan Weß, Klaus-Dieter Althoff, and Michael M. Richter, editors. Topics in case-based reasoning: first European workshop, EWCBR ’93, Kaiserslautern, Germany, November 1–5, 1993: selected papers, volume 837 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58330-0 (Berlin), 0-387-58330-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E93 1993.

Gero:1994:AID

[322] John S. Gero and Fay Sudweeks, editors.

Artificial Intelligence in Design ’94: Proceedings of the Third International Conference on Artificial Intelligence in Design, held August 1994, in Lausanne, Switzerland. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2994-5. LCCN TA174 .A7932 1994.

Hovestadt:1994:ADB

[323] Ludger Hovestadt. A4 — Digitales Bauen: Ein Modell für die weitgehende Computerunterstützung von Entwurf, Konstruktion und Betrieb von Gebäuden. Number 120 in Fortschrittsberichte VDI, Reihe 20, Rechnerunterstützte Verfahren. VDI-Verlag, Düsseldorf, 1994. ISBN ???? LCCN ????

Haton:1995:ACR

[324] Jean-Paul Haton, Mark Keane, and Michel Manago, editors. Advances in case-based reasoning: Second European Workshop, EWCBR-94, Chantilly, France, November 1994: selected papers, volume 984 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60364-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E935 1994.